### v2

# xLSTM-Mixer: Multivariate Time Series Forecasting by Mixing via Scalar Memories

Abstract
Time series data is prevalent across numerous fields, necessitating the development of robust and accurate forecasting models. Capturing patterns both within and between temporal and multivariate components is crucial for reliable predictions. We introduce xLSTM-Mixer, a model designed to effectively integrate temporal sequences, joint time-variate information, and multiple perspectives for robust forecasting. Our approach begins with a linear forecast shared across variates, which is then refined by xLSTM blocks. They serve as key elements for modeling the complex dynamics of challenging time series data. xLSTM-Mixer ultimately reconciles two distinct views to produce the final forecast. Our extensive evaluations demonstrate its superior long-term forecasting performance compared to recent state-of-the-art methods. A thorough model analysis provides further insights into its key components and confirms its robustness and effectiveness. This work contributes to the resurgence of recurrent models in time series forecasting.
1 Introduction
Time series are an essential data modality ubiquitous in many critical fields of application, such as medicine (Hosseini et al., 2021), manufacturing (Essien & Giannetti, 2020), logistics (Seyedan & Mafakheri, 2020), traffic management (Lippi et al., 2013), finance (Lin et al., 2012), audio processing (Latif et al., 2023), and weather modeling (Lam et al., 2023). While significant progress in time series forecasting has been made over the decades, the field is still far from being solved. The regular appearance of yet better models and improved combinations of existing approaches exemplifies this. Further increasing the forecast quality obtained from machine learning models promises a manifold of improvements, such as higher efficiency in manufacturing and transportation as well as more accurate medical treatments.
Historically, recurrent neural networks (RNNs) and their powerful successors were natural choices for deep learning-based time series forecasting (Hochreiter & Schmidhuber, 1997; Cho et al., 2014). Today, large Transformers (Vaswani et al., 2017) are applied extensively to time series tasks, including forecasting. Many improvements to the vanilla architecture have since been proposed, including patching (Nie et al., 2023), decompositions (Zeng et al., 2023), and tokenization inversions (Liu et al., 2023). However, some of their limitations are yet to be lifted. For instance, they typically require large datasets to train successfully, restricting their use to only a subset of conceivable applications. Furthermore, they are inefficient when applied to long sequences due to the cost of the attention mechanism being quadratic in the number of variates and time steps, depending on the specific choice of tokenization. Therefore, recurrent and state space models (SSMs) (Patro & Agneeswaran, 2024) are experiencing a resurgence of interest in overcoming such limitations. Specifically, Beck et al. (2024) revisited recurrent models by borrowing insights gained from Transformers applied to many domains, specifically to natural language processing. They propose Extended Long Short-Term Memory (xLSTM) models as a viable alternative to current sequence models.
We propose xLSTM-Mixer [^1] https://github.com/mauricekraus/xLSTM-Mixer, a new state-of-the-art method for time series forecasting using recurrent deep learning methods. Specifically, we augment the highly expressive xLSTM architecture with carefully crafted time, variate, and multi-view mixing. These operations regularize the training and limit the model parameters by weight-sharing, effectively improving the learning of features necessary for accurate forecasting. xLSTM-Mixer initially computes a channel-independent linear forecast shared over the variates. It is then up-projected to a higher hidden dimension and subsequently refined by an xLSTM stack. It performs multi-view forecasting by producing a forecast from the original and reversed up-projected embedding. The powerful xLSTM cells thereby jointly mix time and variate information to capture complex patterns from the data. Both forecasts are eventually reconciled by a learned linear projection into the final prediction, again by mixing time. An overview of our method is shown in Figure 1.
Figure 1: The xLSTM-Mixer architecture consists of three stages: (1) An initial NLinear forecast assuming channel independence and performing time mixing; (2) subsequent joint mixing, which mixes variate and time information through crucial applications of sLSTM blocks in two views; and (3) a final view mixing, where the two latent forecast views are reconciled into a coherent final forecast.
Overall, we make the following contributions:
  1. (i)
    We investigate time and variate mixing in the context of recurrent models and propose a joint multistage approach that is highly effective for multivariate time series forecasting. We argue that marching over the variates instead of the temporal axis yields better results if suitably combined with temporal mixing.
  2. (ii)
    We propose xLSTM-Mixer, a state-of-the-art method for time series forecasting using recurrent deep learning methods.
  3. (iii)
    We extensively compare xLSTM-Mixer with existing methods for multivariate long-term time series forecasting and perform in-depth model analyses. The experiments demonstrate that xLSTM-Mixer consistently achieves state-of-the-art performance in a wide range of benchmarks.
The following work is structured as follows: In the upcoming Sec. 2, we introduce preliminaries to then motivate and explain xLSTM-Mixer in Sec. 3. We then present comprehensive experiments on its effectiveness and inner workings in Sec. 4. We finally contextualize the findings within the related work in Sec. 5 and close with a conclusion and outlook in Sec. 6.
2 Background
After introducing the notation used throughout this work, we review xLSTM blocks and discuss leveraging channel mixing or their independence in time series models.
2.1 Notation
In multivariate time series forecasting, the model is presented with a time series $\bm{X}=\left(\bm{x}_{1},\dots,\bm{x}_{T}\right)\in\mathbb{R}^{V\times T}$ consisting of $T$ time steps with $V$ variates each. Given this context, the forecaster shall predict the future values $\bm{Y}=\left(\bm{x}_{T+1},\dots,\bm{x}_{T+H}\right)\in\mathbb{R}^{V\times H}$ up to a horizon $H$. A variate (also called a channel) can be any scalar measurement, such as the occupancy of a road or the oil temperature in a power plant. The measurements are assumed to be carried out jointly, such that the $T+H$ time steps reflect a regularly sampled multivariate signal. A time series dataset consists of $N$ such pairs $\left\{\left(\bm{X}^{(i)},\bm{Y}^{(i)}\right)\right\}_{i\in\{1,\dots,N\}}$ divided into train, validation, and test portions.
2.2 Extended Long Short-Term Memory (xLSTM)
Beck et al. (2024) propose xLSTM architectures consisting of two building blocks, namely the sLSTM and mLSTM modules. To harness the full expressivity of xLSTMs within each step and across the computation sequence, we employ a stack of sLSTM blocks without any mLSTM blocks. The latter are less suited for joint mixing due to their independent treatment of the sequence elements, making it impossible to learn any relationships between them directly. We will continue by recalling the construction of sLSTM cells.
The standard LSTM architecture of Hochreiter & Schmidhuber (1997) involves updating the cell state $\mathbf{c}_{t}$ through a combination of input, forget, and output gates, which regulate the flow of information across tokens. sLSTM blocks enhance this by incorporating exponential gating and memory mixing (Greff et al., 2017) to handle complex temporal and cross-variate dependencies more effectively. The sLSTM updates the cell $\bm{c}_{t}$ and hidden state $\bm{h}_{t}$ using three gates as follows:
$\bm{c}_{t}$ $=\bm{f}_{t}\odot c_{t-1}+\bm{i}_{t}\odot\bm{z}_{t}$ cell state (1)
$\bm{n}_{t}$ $=\bm{f}_{t}\cdot\bm{n}_{t-1}+\bm{i}_{t}$ normalizer state (2)
$\bm{h}_{t}$ $=\bm{o}_{t}\odot\bm{c}_{t}\odot\bm{n}^{-1}_{t}$ hidden state (3)
$\bm{z}_{t}$ $=\tanh\bigl{(}\bm{W}_{z}\bm{x}_{t}+\bm{R}_{z}h_{t-1}+\bm{b}_{z}\bigr{)}$ cell input (4)
$\bm{i}_{t}$ $=\exp\bigl{(}\bm{\tilde{i}}_{t}-\bm{m}_{t}\bigr{)}$ $\bm{\tilde{i}}_{t}$ $=\bm{W}_{i}\bm{x}_{t}+\bm{R}_{i}\bm{h}_{t-1}+\bm{b}_{i}$ input gate (5)
$\bm{f}_{t}$ $=\exp\bigl{(}\bm{\tilde{f}}_{t}+\bm{m}_{t-1}-\bm{m}_{t}\bigr{)}$ $\bm{\tilde{f}}_{t}$ $=\bm{W}_{f}\bm{x}_{t}+\bm{R}_{f}\bm{h}_{t-1}+\bm{b}_{f}$ forget gate (6)
$\bm{o}_{t}$ $=\sigma\bigl{(}\bm{W}_{o}\bm{x}_{t}+\bm{R}_{o}\bm{h}_{t-1}+\bm{b}_{o}\bigr{)}$ output gate (7)
$\bm{m}_{t}$ $=\max\bigl{(}\bm{\tilde{f}}_{t}+\bm{m}_{t-1},\bm{\tilde{i}}_{t}\bigr{)}$ stabilizer state (8)
In this setup, the matrices $\bm{W}_{z},\bm{W}_{i},\bm{W}_{f},$ and $\bm{W}_{o}$ are input weights mapping the input token $\bm{x}_{t}$ to the cell input $\bm{z}_{t}$, input gate, forget gate, and output gate, respectively. The states $\bm{n}_{t}$ and $\bm{m}_{t}$ serve as necessary normalization and training stabilization, respectively.
As Beck et al. have shown, it is beneficial to restrict the memory mixing performed by the recurrent weight matrices $\bm{R}_{z},\bm{R}_{i},\bm{R}_{f},$ and $\bm{R}_{o}$ to individual heads, inspired by the multi-head setup of Transformers (Zeng et al., 2023), yet more restricted and therefore more efficient to compute. In particular, each token gets broken up into separate pieces, where the input weights $\bm{W}_{z,i,f,o}$ act across all of them, but the recurrence matrices $\bm{R}_{z,i,f,o}$ are implemented as block-diagonals and therefore only act within each piece. This permits specialization of the individual heads to patterns specific to the respective section of the tokens and empirically does not sacrifice expressivity.
2.3 Channel Independence and Mixing in Time Series Models
Multiple works have investigated whether it is beneficial to learn representations of the time and variate dimensions jointly or separately. Intuitively, because joint mixing is strictly more expressive, one might think it should always be preferred. It is indeed used by many methods such as Temporal Convolutional Networks (TCN) (Lea et al., 2016), N-BEATS (Oreshkin et al., 2019), N-HiTS (Challu et al., 2023), and many Transformers (Vaswani et al., 2017), including Temporal Fusion Transformer (TFT) (Lim et al., 2021), Autoformer (Wu et al., 2021), and FEDFormer (Zhou et al., 2022). However, treating slices of the input data independently assumes an invariance to temporal or variate positions and serves as a strong regularization against overfitting, reminiscent of kernels in CNNs. Prominent models implementing some aspects of channel independence in multivariate time series forecasting are PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2023). TiDE (Das et al., 2023), on the other hand, contains a time-step shared feature projection and temporal decoder but treats variates jointly. As Tolstikhin et al. (2021) have shown with MLP-Mixer, interleaving mixing of all channels in each token and all tokens per channel does not empirically sacrifice any expressivity and instead improves performance. This idea has since been applied to time series, too, namely in architectures such as TimeMixer (Chen et al., 2023c) and TSMixer (Chen et al., 2023c), and is one of the key components of our method xLSTM-Mixer.
3 xLSTM-Mixer
Now we have everything at hand to introduce xLSTM-Mixer as depicted in Fig. 1. It carefully integrates three key components: (1) an initial linear forecast with time mixing, (2) joint mixing using powerful sLSTM modules, and (3) an eventual combination of two views by a final fully connected layer. The transposing steps between the key components enable capturing complex temporal and intra-variate patterns while facilitating easy trainability and limiting parameter counts. The sLSTM block, in particular, can learn intricate non-linear relationships hidden within the data along both the time and variate dimensions. The xLSTM-Mixer architecture is furthermore equipped with normalization layers and skip connections to improve training stability and overall effectiveness.
3.1 Key Component 1: Normalization and Initial Linear Forecast
Normalization has become an essential ingredient of modern deep learning architectures (Huang et al., 2023). For time series in particular, reversible instance norm (RevIN) (Kim et al., 2022) is a general recipe for improving forecasting performance, where each time series instance is normalized by its mean and variance and furthermore scaled and offset by learnable scalars $\gamma$ and $\beta$:
$\bm{x}_{t}^{\text{norm}}=\operatorname{RevIN}(\bm{x}_{t})=\gamma\left(\frac{\bm{x}_{t}-\mathbb{E}\left[\bm{x}\right]}{\sqrt{\mathrm{Var}\left[\bm{x}\right]}+\epsilon}\right)+\beta.$
We apply it as part of xLSTM-Mixer, and at the end of the entire pipeline, we invert the RevIN operation to obtain the final prediction. In the case of xLSTM-Mixer, the typical skip connections found in mixer acrchitectures (Tolstikhin et al., 2021; Chen et al., 2023c) are taken up by RevIN, the normalization in the NLinear forecast explained shortly, and the integral skip connections within each sLSTM block.
It has been shown previously that simple linear models equipped with appropriate normalization schemes are, already by themselves, decent long-term forecasters (Zeng et al., 2023; Li et al., 2023). Our observations confirm this finding. Therefore, we first process each variate separately by an NLinear model by computing:
$\bm{x}^{\text{initial}}=\operatorname{NLinear}(\bm{x}^{\text{norm}})=\operatorname{FC}\left(\bm{x}_{1:T}^{\text{norm}}-x_{T}^{\text{norm}}\right)+x_{T}^{\text{norm}}\,,$
where $\operatorname{FC}(\cdot)$ denotes a fully-connected linear layer with bias term. Sharing this model across variates limits parameter counts, and the weight-tying serves as a useful regularization. The quality of this initial forecast will be investigated in Sec. 4.1 and 4.2.
3.2 Key Component 2: sLSTM Refinement
While the NLinear forecast $\bm{x}^{\text{initial}}$ captures the basic patterns between the historic and future time steps, its quality alone is insufficient for today’s challenging time series datasets. We, therefore, refine it using powerful sLSTM blocks. As a first step, it is crucial to increase the embedding dimension of the data to provide enough latent dimensions $D$ for the sLSTM cells: $\bm{x}^{\text{up}}=\operatorname{FC}^{\text{up}}\left(\bm{x}^{\text{initial}}\right)$. This pre-up-projection is similar to what is commonly performed in SSMs (Beck et al., 2024). We weight-share $\operatorname{FC}^{\text{up}}$ across variates to perform time-mixing similar to the initial forecast. Note that this step does not maintain the temporal ordering within the embedding token dimensions, as was the case up until this step, and instead embeds it into a higher latent dimension.
The stack of $M$ sLSTM blocks $\mathcal{S}(\cdot)$ transforms $\bm{x}^{\text{up}}$ as defined in Eq. 1 to 8. The recurrent model strides over the data in variate order, i.e., where each token represents all time steps from a single variate as in the work of Liu et al. (2023). The sLSTM blocks learn intricate non-linear relationships hidden within the data along both the time and variate dimensions. The mixing of the hidden state is still limited to blocks of consecutive dimensions, aiding efficient learning and inference while allowing for effective cross-variate interaction during the recurrent processing. Striding over variates has the benefit of linear runtime scaling in the number of variates at a constant number of parameters. It, however, comes at the cost of possibly fixing a suboptimal order of variates. While this is empirically not a significant limitation, we leave investigations into how to find a suitable ordering for future work. In addition to a large embedding dim, we observed a high number of heads being crucial for effective forecasting.
The sLSTM cells’ first hidden state $\bm{h}_{t-1}$ must be initialized before each sequence of tokens can be processed. Extending the initial description of these blocks, we propose learning a single initial embedding token $\bm{\eta}\in\mathbb{R}^{D}$ that gets prepended to each encoded time series $\bm{x}^{\text{up}}$. These initial embeddings draw from recent advances in Large Language Models, where learnable "soft prompt" tokens are used to condition models and improve their ability to generate coherent outputs (Lester et al., 2021; Li & Liang, 2021; Chen et al., 2023a; b). Recent research has extended the application of soft prompts to LLM-based time series forecasting (Cao et al., 2023; Sun et al., 2024), emphasizing their adaptability and effectiveness in improving model performance across modalities. These tokens enable greater flexibility and conditioning, allowing the model to adapt its initial memory representation to specific dataset characteristics and to dynamically interact with the time and variate data. Soft prompts can be readily optimized through back-propagation with very little overhead.
3.3 Key Component 3: Multi-View Mixing
To further regularize the training of the sLSTM as with the linear projections, we compute forecasts from the original embedding $\bm{x}^{\text{up}}$ as well as the reversed embedding $\bm{\widehat{x}}^{\text{up}}$, where the order of the latent dimensions including the representation of $\bm{\eta}$ is inverted. Learning forecasts $\bm{y}^{\prime}$ and $\bm{y}^{\prime\prime}$ for both views while sharing weights helps learn better representations. Such multi-task learning settings are known to benefit training (Zhang & Yang, 2022). The final forecast is obtained by a linear projection $\operatorname{FC}^{\text{view}}$ of the two concatenated forecasts, again per-variate. Specifically, we compute:
[…]
The final forecast is obtained after de-normalizing the reconciled forecasts as $\bm{y}=\operatorname{RevIN}^{-1}(\bm{y}^{\text{norm}})$.
4 Experimental Evaluation
Our intention here is to evaluate the forecasting capabilities of xLSTM-Mixer, aiming to provide comprehensive insights into its performance. To this end, we conducted a series of experiments with the primary focus on long-term forecasting, following the work of Das et al. (2023) and Chen et al. (2023c). An evaluation of xLSTM-Mixer’s competitiveness in short-term forecasting on the PEMS dataset is provided in Sec. A.2. Additionally, we perform an extensive model analysis consisting of an ablation study to identify the contributions of individual components of xLSTM-Mixer, followed by an inspection of the initial embedding tokens, a hyperparameter sensitivity analysis, and an investigation into its robustness.
Table 1: The long-term forecasting benchmark datasets and their key properties.
Dataset Source Domain Horizons Sampling #Variates
Weather Zhou et al. (2021) Weather 96–720 10 min 21
Electricity Zhou et al. (2021) Power Usage 96–720 1 hour 321
Traffic Wu et al. (2021) Traffic Load 96–720 1 hour 862
ETT Zhou et al. (2021) Power Production 96–720 15&60 min 7
Datasets.   We generally follow the established benchmark procedure of Wu et al. (2021) and Zhou et al. (2021) for best backward and future comparability. The datasets we thus used are summarized in Tab. 1.
Training.   We follow standard practice in the forecasting literature by evaluating long-term forecasts using the mean squared error (MSE) and the mean absolute error (MAE). Based on our experiments, we used the MAE as the training loss function since it yielded the best results. The datasets were standardized for consistency across features. Further details on hyperparameter selection, metrics, and the implementation can be found in Sec. A.1.
Baseline Models.   We compare xLSTM-Mixer to the recurrent models xLSTMTime (Alharthi & Mahmood, 2024) and LSTM (Hochreiter & Schmidhuber, 1997); multi-perceptron (MLP) based models TimeMixer (Wang et al., 2024a), TSMixer (Chen et al., 2023c), DLinear (Zeng et al., 2023), and TiDE (Das et al., 2023); the Transformers PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2023), FEDFormer (Zhou et al., 2022), and Autoformer (Wu et al., 2021); and the convolutional architectures MICN (Wang et al., 2022) and TimesNet (Wu et al., 2022).
4.1 Long-Term Time Series Forecasting
Table 2: xLSTM-Mixer is effective in long-term forecasting. The results are averaged from 4 different prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction. The best result for each dataset is highlighted in bold red, while the second-best result is blue and underlined. Wins for each model out of all 28 settings are shown at the bottom.
Models Recurrent MLP Transformer Convolutional
xLSTM-Mixer
xLSTMTime
LSTM
TimeMixer
TSMixer
DLinear
TiDE
PatchTST
iTransformer
FEDFormer
Autoformer
MICN
TimesNet
1997a
Dataset
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Weather
0.219
0.250
0.222 0.255
0.444
0.454
0.222
0.262
0.225
0.264
0.246
0.300
0.236
0.282
0.241
0.264
0.258
0.278
0.309
0.360
0.338
0.382
0.242
0.299
0.259
0.287
Electricity
0.153
0.245
0.157
0.250
0.559
0.549
0.156 0.246
0.160
0.256
0.166
0.264
0.159
0.257
0.159
0.253
0.178
0.270
0.214
0.321
0.227
0.338
0.186
0.295
0.192
0.295
Traffic
0.392
0.253
0.391
0.261
1.011
0.541
0.387
0.262
0.408
0.284
0.434
0.295
0.356
0.261
0.391
0.264
0.428
0.282
0.609
0.376
0.628
0.379
0.541
0.315
0.620
0.336
ETTh1
0.397
0.420
0.408
0.428
1.198
0.821
0.411
0.423
0.412
0.428
0.423
0.437
0.419
0.430
0.413
0.434
0.454
0.448
0.440
0.460
0.496
0.487
0.558
0.535
0.458
0.450
ETTh2
0.340
0.382
0.346
0.386
3.095
1.352
0.316
0.384
0.355
0.401
0.431
0.447
0.345
0.394
0.324
0.381
0.383
0.407
0.433
0.447
0.453
0.462
0.588
0.525
0.414
0.427
ETTm1
0.339
0.366
0.347 0.372
1.142
0.782
0.348
0.375
0.347
0.375
0.357
0.379
0.355
0.378
0.353
0.382
0.407
0.410
0.448
0.452
0.588
0.517
0.392
0.413
0.400
0.406
ETTm2
0.248
0.307
0.254
0.310
2.395
1.177
0.256
0.315
0.267
0.322
0.267
0.332
0.249
0.312
0.256
0.317
0.288
0.332
0.304
0.349
0.324
0.368
0.328
0.382
0.291
0.333
Wins
18
23
1
2
0
0
3
5
0
0
0
0
5
1
2
1
0
0
0
0
0
0
0
0
0
0
  • a
    Taken from Wu et al. (2022).
Figure 2: Example Forecasts Across Models and Datasets. This figure shows example forecasts on the Weather and ETTm1 datasets for multiple models. The blue lines represent the ground truth target, and the orange lines show the predictions. The first panel illustrates the forecast from xLSTM-Mixer, while the second panel shows the forecast extracted before the up-projection step, highlighting the effectiveness of our added components. Comparisons with xLSTMTime, TimeMixer, and PatchTST confirm the favorable performance of xLSTM-Mixer relative to these baseline models.
We present the performance of xLSTM-Mixer compared to prior models in Tab. 2. As shown, xLSTM-Mixer consistently delivers highly accurate forecasts across a wide range of datasets. It achieves the best results in 18 out of 28 cases for MSE and 22 out of 28 cases for MAE, demonstrating its superior performance in long-term forecasting. In particular, xLSTM-Mixer exhibits exceptional forecasting accuracy, as evidenced particularly by its strong MAE performance across all datasets. Notably, on Weather, xLSTM-Mixer reduces the MAE by 2% compared to xLSTMTime and 4.6% compared to TimeMixer. Similarly, for ETTm1, xLSTM-Mixer outperforms TimeMixer by 2.4% in MAE and shows a strong competitive edge over xLSTMTime. Although xLSTM-Mixer performs slightly less well on the Traffic and ETTh2 datasets, where it encounters challenges with handling outliers, it remains highly competitive and outperforms the majority of baseline models. This suggests that despite these few cases, xLSTM-Mixer can consistently deliver state-of-the-art performance in long-term forecasting. A qualitative inspection of several baseline models, including the initial forecast extracted before the sLSTM refinement, is shown in Fig. 2. In this comparison, the lookback window and forecasting horizon are both fixed at 96.
4.2 Model Analysis
Ablation Study.
To assess the contributions of each component in xLSTM-Mixer to its strong overall forecast performance, we conducted an extensive ablation study with the results listed in Tab. 3. Each configuration represents a different combination of the four key components: mixing time with NLinear, using sLSTM blocks, learning an initial embedding token, and multi-view mixing. We evaluated the performance using the MSE and MAE across the prediction lengths $\{96,192,336,720\}$.
Table 3: Ablation Study on the Weather and ETTm1 Datasets. The MSE (↓) and MAE (↓) are reported for each numbered configuration across all four forecast horizons. The best results are highlighted in bold red, while the second-best results are blue and underlined.
#1 (full)
#2
#3
#4
#5
#6
#7
#8
#9
#10
Mix Time
sLSTM
Variates
Time
Variates
Variates
Variates
None
Variates
Variates
Variates
Variates
Init. Token
Mix View
Horizon
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Weather
96
0.143
0.184
0.148
0.194
0.145
0.186
0.144
0.185
0.144
0.186
0.173
0.223
0.149
0.193
0.151
0.195
0.149
0.192
0.152
0.195
192
0.186
0.226
0.196
0.239
0.188
0.228
0.186
0.226
0.188
0.228
0.219
0.257
0.192
0.233
0.192
0.234
0.191
0.234
0.193
0.236
336
0.237
0.266
0.252
0.281
0.239
0.267
0.241
0.270
0.242
0.270
0.261
0.288
0.240
0.271
0.242
0.273
0.242
0.273
0.244
0.274
720
0.310
0.324
0.315
0.328
0.310
0.324
0.309
0.323
0.309
0.323
0.320
0.334
0.320
0.329
0.319
0.329
0.322
0.330
0.319
0.328
ETTm1
96
0.275
0.328
0.298
0.348
0.277
0.329
0.278
0.331
0.279
0.333
0.295
0.338
0.282
0.339
0.285
0.341
0.281
0.337
0.284
0.339
192
0.319
0.354
0.337
0.369
0.321
0.354
0.321
0.356
0.322
0.358
0.329
0.357
0.329
0.364
0.330
0.365
0.337
0.367
0.335
0.366
336
0.353
0.374
0.368
0.388
0.354
0.375
0.355
0.377
0.357
0.379
0.359
0.376
0.367
0.385
0.367
0.385
0.366
0.384
0.366
0.385
720
0.409
0.407
0.420
0.416
0.411
0.408
0.413
0.411
0.414
0.411
0.412
0.407
0.422
0.412
0.422
0.413
0.417
0.410
0.418
0.411
The full version of xLSTM-Mixer (#1), which integrates all components, achieves the best performance overall. However, we also observe that some configurations of xLSTM-Mixer, which exclude specific components, remain competitive. For instance, #3, which excludes the initial embedding token, still performs reasonably well. This suggests that while it contributes positively to the overall performance, the model can sometimes still achieve competitive results without it. In general, removing any specific component leads to a performance drop. For example, removing the time mixing (#7) increases the MAE by 3.4% on ETTm1 at length 96 or 2.8% at length 192, highlighting its critical role in capturing intratemporal dependencies. When we now omit everything except for time mixing on Weather at 192, we suffer a 13.7% performance decrease. In summary, the ablation study confirms that all components of xLSTM-Mixer contribute to its effectiveness, with the full configuration yielding the best results. Furthermore, we identified the sLSTM blocks and time-mixing as critical components for ensuring high accuracy across datasets and prediction lengths.
Figure 3: Initial tokens capture dataset characteristics. The plot illustrates the learned tokens across multiple datasets and prediction lengths. The lookback length is set to 96 for all evaluations. For clarity and the high noise levels in ETTm1 and ETTh2, we only show three seeds for Weather.
Initial Token Embedding.
We qualitatively inspect decodings of the initial embedding tokens $\bm{\eta}$ on multiple datasets to further understand and interpret the initializations learned by xLSTM-Mixer. $\bm{\eta}$ are decoded to a forecast $\bm{y}$ by transforming them through the sLSTM stack $\mathcal{S}$ and applying multi-view mixing. The resulting output of $\operatorname{FC}^{\text{view}}$ can then be interpreted as the conditioning forecast used to initialize the sLSTM blocks. Fig. 3 shows the dataset-specific patterns the initial embedding tokens have learned on Weather, ETTm1, and ETTh2 for various prediction horizons. With increasing prediction horizons, we observe longer spans of time, eventually revealing underlying seasonal patterns and respective dataset dynamics.
Figure 4: Sensitivity analysis of the xLSTM hidden dimension $D$. The shaded area indicates one standard deviation from the mean MSE (↓). Increasing the up-projection dimension becomes beneficial with increasing prediction length.
Sensitivity to xLSTM Hidden
Dimension.
In Fig. 4, we visualize the performance of xLSTM-Mixer on the Electricity dataset with increasing sLSTM embedding (hidden) dimension realized by $\operatorname{FC}^{\text{up}}$. The results indicate that larger hidden dimensions consistently enhance the model’s performance, particularly for longer forecast horizons. This suggests that a larger embedding dimension enables xLSTM-Mixer to capture better the higher complexity of the time series data over extended horizons, leading to improved forecasting accuracy.
Robustness to Lookback Length.
Fig. 5 illustrates the performance of xLSTM-Mixer across varying lookback lengths and prediction horizons. We observe that xLSTM-Mixer can effectively utilize longer lookback windows than the baselines, especially when compared to transformer-based models. This advantage stems from xLSTM-Mixer’s avoidance of self-attention, allowing it to handle extended lookback lengths efficiently. Additionally, xLSTM-Mixer demonstrates stable and consistent performance with low variance. These results confirm that increasing the lookback length improves forecasting accuracy and enhances robustness, particularly for longer prediction horizons.
Figure 5: Increasing the lookback window increases forecasting performance, with xLSTM-Mixer virtually always providing the best results. Shows the mean MSE (↓) and std. dev. on ETTm1.
5 Related Work
Time Series Forecasting.
A long line of machine learning research led from early statistical methods like ARIMA (Box & Jenkins, 1976) to contemporary models based on deep learning, where four architectural families take center stage: The ones based on recurrence, convolutions, Multilayer Perceptrons (MLPs), and Transformers. While all of them are used by practitioners today, the research focus is gradually shifting over time. Initially, the naturally sequential recurrent models such as Long Short-Term Memory (LSTM) (Hochreiter & Schmidhuber, 1997) and Gated Recurrent Units (GRUs) (Cho et al., 2014) were used for time series analysis. Their main benefits are the high inference efficiency and arbitrary input and output lengths due to their autoregressive nature. While their effectiveness has historically been constrained by a limited ability to capture long-range dependencies, active research continues to alleviate these limitations (Salinas et al., 2020), including the xLSTM architecture presented in Sec. 2 (Beck et al., 2024; Alharthi & Mahmood, 2024). Similarly efficient as RNNs, yet more restricted in their output length, are the location-invariant CNNs (Li et al., 2022; Lara-Benítez et al., 2021), such as TCN (Lea et al., 2016), TimesNet (Wu et al., 2022), and MICN (Wang et al., 2022). Recently, some MLP-based architectures have also shown good success, including the simplistic DLinear and NLinear models (Zeng et al., 2023), the encoder-decoder architecture of TiDE (Das et al., 2023), the mixing architectures TimeMixer (Wang et al., 2024a) and TSMixer (Chen et al., 2023c), as well as the hierarchical N-BEATS Oreshkin et al. (2019) and N-HiTS (Challu et al., 2023) models. Finally, a lot of models have been proposed based on Transformers (Vaswani et al., 2017), such as Autoformer (Wu et al., 2021), TFT (Lim et al., 2021), FEDFormer (Zhou et al., 2022), PatchTST (Nie et al., 2023), and iTransformer (Liu et al., 2023).
xLSTM Models for Time Series.
Some initial experiments of applying xLSTMs (Beck et al., 2024) to time series were already performed by Alharthi & Mahmood (2024) with their proposed xLSTMTime model. While it showed promising forecasting performance, these initial soundings did not surpass stronger recent models such as TimeMixer (Wang et al., 2024a) on multivariate benchmarks, and the reported performance is challenging to reproduce. We ensure that our method xLSTM-Mixer is well suited as a foundation for further research by providing extensive model analysis, including an ablation study with ten variants, and ensuring that results are readily reproducible. Our methodology draws from xLSTMTime yet improves on it by several key components. Most importantly, our novel multi-view mixing consistently enhances forecasting performance. Furthermore, we find the trend-seasonality decomposition to be redundant and a simple NLinear normalization scheme (Zeng et al., 2023) to suffice.
6 Conclusion
In this work, we introduced xLSTM-Mixer, a method that combines a linear forecast with further refinement using xLSTM blocks. Our architecture effectively integrates time, joint, and view mixing to capture complex dependencies. In long-term forecasting, xLSTM-Mixer consistently achieved state-of-the-art performance, outperforming previous methods in 41 out of 56 cases. Furthermore, our detailed model analysis provided valuable insights into the contribution of each component and demonstrated its robustness to varying hyperparameter settings.
While xLSTM-Mixer has shown extraordinary performance in long-term forecasting, it should be noted that due to the transpose of the input, i.e., processing the variates as sequence elements, the number of variates may limit the overall performance. To overcome this, we plan to explore how different variate orderings influence performance and whether incorporating more than two views could lead to further improvements. This study focused on long-term forecasting, yet extending xLSTM-Mixer to tasks such as short-term forecasting, time series classification, or imputation offers promising directions for future research.
Ethics Statement
Our research advances machine learning by enhancing the capabilities of long-term forecasting in time series models, significantly improving both accuracy and efficiency. By developing xLSTM-Mixer, we introduce a robust framework that can be applied across various industries, including finance, healthcare, energy, and logistics. The improved forecasting accuracy enables better decision-making in critical areas, such as optimizing resource allocation, predicting market trends, and managing risk.
However, we also recognize the potential risks associated with the misuse of these advanced models. Time series forecasting models could be leveraged for malicious purposes, especially when applied at scale. For example, in the financial sector, adversarial agents might manipulate forecasts to create market instability. In political or social contexts, these models could be exploited to predict and influence public opinion or destabilize economies. Additionally, the application of these models in sensitive domains like healthcare and security may lead to unintended consequences if not carefully regulated and ethically deployed.
Therefore, it is essential that the use of xLSTM-Mixer, like all machine learning technologies, is guided by responsible practices and ethical considerations. We encourage stakeholders to adopt rigorous evaluation processes to ensure fairness, transparency, and accountability in its deployment, and to remain vigilant to the broader societal implications of time series forecasting technologies.
Reproducibility Statement
All implementation details, including dataset descriptions, metric calculations, and experiment configurations, are provided in Sec. 4 and Sec. A.1. We make sure to exclusively use openly available software and datasets and provide the source code for full reproducibility at
https://github.com/mauricekraus/xLSTM-Mixer.
Acknowledgments
This work received funding from the EU project EXPLAIN, under the Federal Ministry of Education and Research (BMBF) (grant 01—S22030D). Furthermore, it was funded by the ACATIS Investment KVG mbH project “Temporal Machine Learning for Long-Term Value Investing” and the BMBF KompAKI project within the “The Future of Value Creation – Research on Production, Services and Work” program (funding number 02L19C150), managed by the Project Management Agency Karlsruhe (PTKA). The author of Eindhoven University of Technology received support from their Department of Mathematics and Computer Science and the Eindhoven Artificial Intelligence Systems Institute. Furthermore, this work benefited from the HMWK project “The Third Wave of Artificial Intelligence - 3AI”.
Appendix A Appendix
A.1 Implementation Details
Experimental Details
Our codebase is implemented in Python 3.11, leveraging PyTorch 2.4  (Paszke et al., 2019) in combination with Lightning 2.4222https://lightning.ai/pytorch-lightning for model training and optimization. We used the custom CUDA implementation of Beck et al. (2024) for sLSTM [^3] https://github.com/NX-AI/xlstm which relies on the NVIDIA Compute Capability 8.0 or higher. Thus, our experiments were conducted on a single NVIDIA A100 80GB GPU. The majority of our baseline implementations, along with data loading and preprocessing steps, are adapted from the Time-Series-Library [^4] https://github.com/thuml/Time-Series-Library of Wang et al. (2024b). Additionally, for xLSTMTime we used code based on the official repository [^5] https://github.com/muslehal/xLSTMTime of Alharthi & Mahmood (2024).
Training and Hyperparameters
We optimized xLSTM-Mixer for up to 60 epochs with a cosine-annealing scheduler with the Adam optimizer (Kingma & Ba, 2017), using $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and no weight decay. Hyperparameter tuning was conducted using Optuna (Akiba et al., 2019) with the choices provided in Tab. 4. We optimized for the L1 forecast error, also known as the Mean Absolute Error (MAE). To further stabilize the training process, gradient clipping with a maximum norm of $1.0$ was applied. All experiments were run with three different random seeds {2021, 2022, 2023}.
Table 4: Hyperparameters and their choices.
Hyperparameter Choices
Batch size {16, 32, 64, 128, 256, 512}
Initial learning rate $1\cdot 10^{-2}$, $3\cdot 10^{-3}$, $1\cdot 10^{-3}$, $5\cdot 10^{-4}$, $2\cdot 10^{-4}$, $1\cdot 10^{-4}$}
Scheduler warmup steps {5, 10, 15}
Lookback length $T$96, 256, 512, 768, 1024, 2048}
Embedding dimension $D$32, 64, 128, 256, 512, 768, 1024}
sLSTM conv. kernel width {disabled, 2, 4}
sLSTM dropout rate {0.1, 0.25}
# sLSTM blocks in $\mathcal{S}$1, 2, 3, 4}
# sLSTM heads {4, 8, 16, 32}
Metrics
We follow common practice in the literature (Wu et al., 2021; Wang et al., 2024a) for maximum comparability and, therefore, evaluate long-term forecasting of all models on the mean absolute error (MAE), mean squared error (MSE), and for short-term forecasting, using the MAE, root mean squared error (RMSE), and mean absolute percentage error (MAPE). The metrics are averaged over all variates and computed as:
$\operatorname{MAE}(\bm{y},\bm{\hat{y}})$ $=\sum_{i=1}^{H}\left|y_{i}-\hat{y_{i}}\right|$ $\operatorname{MSE}(\bm{y},\bm{\hat{y}})$ $=\sum_{i=1}^{H}(y_{i}-\hat{y_{i}})^{2}$
$\operatorname{RMSE}(\bm{y},\bm{\hat{y}})$ $=\sqrt{\operatorname{MSE}(\bm{y},\bm{\hat{y}})}$ $\operatorname{MAPE}(\bm{y},\bm{\hat{y}})$ $=\frac{100}{H}\sum_{i=1}^{H}\frac{\left|y_{i}-\hat{y}_{i}\right|}{\left|y_{i}\right|+\epsilon},$
where $\bm{y}$ are the targets, $\bm{\hat{y}}$ the predictions, and $\epsilon$ a small constant added for numerical stability.
A.2 Outlook: Short-Term Time Series Forecasting
Having shown superior long-term forecasting accuracies in Sec. 4.1, we also provide an initial exploration of the effectiveness of xLSTM-Mixer to short-term forecasts. To this end, we compare it to applicable baselines on PEMS datasets with input lengths uniformly set to 96 and prediction lengths to 12. The results in Tab. 5 show that the performance of xLSTM-Mixer is competitive with existing methods. We provide the MAE, MAPE, and RMSE as is common practice.
Table 5: Short-term forecasting evaluation of xLSTM-Mixer and baselines on the multivariate PEMS datasets. A lower MAE, MAPE, or RMSE indicates a better prediction. The best result for each dataset is highlighted in bold red, while the second-best result is blue and underlined.
Models Recurrent MLP Transformer Convolutional
xLSTM-Mixer
xLSTMTime
LSTM
TimeMixer
DLinear
PatchTST
FEDFormer
Autoformer
MICN
TimesNet
1997a
PEMS03
MAE
15.71
16.59
18.65
14.63
19.70
18.95
19.00
18.08
15.71
16.41
MAPE
14.92
15.31
17.39
14.54
18.35
17.29
18.57
18.75
15.67
15.17
RMSE
24.82
26.47
31.73
23.28
32.35
30.15
30.05
27.82
24.55
26.72
PEMS08
MAE
16.56
17.44
20.34
15.22
20.26
20.35
20.56
20.47
17.76
19.01
MAPE
10.24
10.58
13.05
9.67
12.09
13.15
12.41
12.27
10.76
11.83
RMSE
26.65
28.13
31.90
24.26
32.38
31.04
32.97
31.52
27.26
30.65
  • a
    Configuration following Wu et al. (2021).
A.3 Full Results for Long-Term Forecasting
Tab. 6 shows the full results for long-term forecasting on all four separate forecast horizons.
Table 6: Full long-term forecasting results for Tab. 2. Avg is averaged from all four prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction. The best result for each dataset is highlighted in bold red, while the second-best result is blue and underlined. Wins for each model are shown at the bottom.
Models Recurrent MLP Transformer Convolutional
xLSTM-Mixer
xLSTMTime
LSTM
TimeMixer
TSMixer
DLinear
TiDE
PatchTST
iTransformer
FEDFormer
Autoformer
MICN
TimesNet
1997a
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Weather
96
0.143
0.184
0.144 0.187
0.369
0.406
0.147
0.197
0.145
0.198
0.176
0.237
0.166
0.222
0.149
0.198
0.174
0.214
0.217
0.296
0.266
0.336
0.161
0.229
0.172
0.220
192
0.186
0.226
0.192
0.236
0.416
0.435
0.189
0.239
0.191
0.242
0.220
0.282
0.209
0.263
0.194
0.241
0.221
0.254
0.276
0.336
0.307
0.367
0.220
0.281
0.219
0.261
336
0.236
0.266
0.237 0.272
0.455
0.454
0.241
0.280
0.242
0.280
0.265
0.319
0.254
0.301
0.306
0.282
0.278
0.296
0.339
0.380
0.359
0.395
0.278
0.331
0.280
0.306
720
0.310
0.323
0.313
0.326
0.535
0.520
0.310
0.330
0.320
0.336
0.323
0.362
0.313
0.340
0.314
0.334
0.358
0.347
0.403
0.428
0.419
0.428
0.311
0.356
0.365
0.359
Avg
0.219
0.250
0.222 0.255
0.444
0.454
0.222
0.262
0.225
0.264
0.246
0.300
0.236
0.282
0.241
0.264
0.258
0.278
0.309
0.360
0.338
0.382
0.242
0.299
0.259
0.287
Electricity
96
0.126
0.218
0.128 0.221
0.375
0.437
0.129
0.224
0.131
0.229
0.140
0.237
0.132
0.229
0.129
0.222
0.148
0.240
0.193
0.308
0.201
0.317
0.164
0.269
0.168
0.272
192
0.144 0.235
0.150
0.243
0.442
0.473
0.140
0.220
0.151
0.246
0.153
0.249
0.147
0.243
0.147
0.240
0.162
0.253
0.201
0.315
0.222
0.334
0.177
0.285
0.184
0.289
336
0.157
0.250
0.166
0.259
0.439
0.473
0.161 0.255
0.161
0.261
0.169
0.267
0.161
0.261
0.163
0.259
0.178
0.269
0.214
0.329
0.231
0.338
0.193
0.304
0.198
0.300
720
0.183
0.276
0.185
0.276
0.980
0.814
0.194
0.287
0.197
0.293
0.203
0.301
0.196
0.294
0.197
0.290
0.225
0.317
0.246
0.355
0.254
0.361
0.212
0.321
0.220
0.320
Avg
0.153
0.245
0.157
0.250
0.559
0.549
0.156 0.246
0.160
0.256
0.166
0.264
0.159
0.257
0.159
0.253
0.178
0.270
0.214
0.321
0.227
0.338
0.186
0.295
0.192
0.295
Traffic
96
0.357
0.236
0.358
0.242
0.843
0.453
0.360
0.249
0.376
0.264
0.410
0.282
0.336
0.253
0.360
0.249
0.395
0.268
0.587
0.366
0.613
0.388
0.519
0.309
0.593
0.321
192
0.377
0.241
0.378
0.253
0.847
0.453
0.375 0.250
0.397
0.277
0.423
0.287
0.346
0.257
0.379
0.256
0.417
0.276
0.604
0.373
0.616
0.382
0.537
0.315
0.617
0.336
336
0.394
0.250
0.392
0.261
0.853
0.455
0.385
0.270
0.413
0.290
0.436
0.296
0.355
0.260
0.392
0.264
0.433
0.283
0.621
0.383
0.622
0.337
0.534
0.313
0.629
0.336
720
0.439
0.283
0.434
0.287
1.500
0.805
0.430 0.281
0.444
0.306
0.466
0.315
0.386
0.273
0.432
0.286
0.467
0.302
0.626
0.382
0.660
0.408
0.577
0.325
0.640
0.350
Avg
0.392
0.253
0.391
0.261
1.011
0.541
0.387
0.262
0.408
0.284
0.434
0.295
0.356
0.261
0.391
0.264
0.428
0.282
0.609
0.376
0.628
0.379
0.541
0.315
0.620
0.336
ETTh1
96
0.359
0.386
0.368
0.395
1.044
0.773
0.361
0.390
0.361
0.392
0.375
0.399
0.375
0.398
0.370
0.400
0.386
0.405
0.376
0.419
0.449
0.459
0.421
0.431
0.384
0.402
192
0.402
0.417
0.401
0.416
1.217
0.832
0.409
0.414
0.404
0.418
0.405
0.416
0.412
0.422
0.413
0.429
0.441
0.436
0.420
0.448
0.500
0.482
0.474
0.487
0.436
0.429
336
0.408
0.429
0.422
0.437
1.259
0.841
0.430
0.429
0.420
0.431
0.439
0.443
0.435
0.433
0.422
0.440
0.487
0.458
0.459
0.465
0.521
0.496
0.569
0.551
0.491
0.469
720
0.419
0.448
0.441
0.465
1.271
0.838
0.445
0.460
0.463
0.472
0.472
0.490
0.454
0.465
0.447
0.468
0.503
0.491
0.506
0.507
0.514
0.512
0.770
0.672
0.521
0.500
Avg
0.397
0.420
0.408
0.428
1.198
0.821
0.411
0.423
0.412
0.428
0.423
0.437
0.419
0.430
0.413
0.434
0.454
0.448
0.440
0.460
0.496
0.487
0.558
0.535
0.458
0.450
ETTh2
96
0.267
0.329
0.273
0.333
2.522
1.278
0.271
0.330
0.274
0.341
0.289
0.353
0.270
0.336
0.274
0.337
0.297
0.349
0.346
0.388
0.358
0.397
0.299
0.364
0.340
0.374
192
0.338
0.375
0.340
0.378
3.312
1.384
0.317
0.402
0.339
0.385
0.383
0.418
0.332
0.380
0.314
0.382
0.380
0.400
0.429
0.439
0.456
0.452
0.441
0.454
0.402
0.414
336
0.367
0.401
0.373
0.403
3.291
1.388
0.332 0.396
0.361
0.406
0.448
0.465
0.360
0.407
0.329
0.384
0.428
0.432
0.496
0.487
0.482
0.486
0.654
0.567
0.452
0.452
720
0.388
0.424
0.398
0.430
3.257
1.357
0.342
0.408
0.445
0.470
0.605
0.551
0.419
0.451
0.379 0.422
0.427
0.445
0.463
0.474
0.515
0.511
0.956
0.716
0.462
0.468
Avg
0.340
0.382
0.346
0.386
3.095
1.352
0.316
0.384
0.355
0.401
0.431
0.447
0.345
0.394
0.324
0.381
0.383
0.407
0.433
0.447
0.453
0.462
0.588
0.525
0.414
0.427
ETTm1
96
0.275
0.328
0.286
0.335
0.863
0.664
0.291
0.340
0.285
0.339
0.299
0.343
0.306
0.349
0.293
0.346
0.334
0.368
0.379
0.419
0.505
0.475
0.316
0.362
0.338
0.375
192
0.319
0.354
0.329
0.361
1.113
0.776
0.327
0.365
0.327
0.365
0.335
0.365
0.335
0.366
0.333
0.370
0.377
0.391
0.426
0.441
0.553
0.496
0.363
0.390
0.374
0.387
336
0.353
0.374
0.358
0.379
1.267
0.832
0.360
0.381
0.356
0.382
0.369
0.386
0.364
0.384
0.369
0.392
0.426
0.420
0.445
0.459
0.621
0.537
0.408
0.426
0.410
0.411
720
0.409
0.407
0.416
0.411
1.324
0.858
0.415
0.417
0.419
0.414
0.425
0.421
0.413
0.413
0.416
0.420
0.491
0.459
0.543
0.490
0.671
0.561
0.481
0.476
0.478
0.450
Avg
0.339
0.366
0.347 0.372
1.142
0.782
0.348
0.375
0.347
0.375
0.357
0.379
0.355
0.378
0.353
0.382
0.407
0.410
0.448
0.452
0.588
0.517
0.392
0.413
0.400
0.406
ETTm2
96
0.157
0.244
0.164
0.250
2.041
1.073
0.164
0.254
0.163
0.252
0.167
0.260
0.161
0.251
0.166
0.256
0.180
0.264
0.203
0.287
0.255
0.339
0.179
0.275
0.187
0.267
192
0.213
0.285
0.218
0.288
2.249
1.112
0.223
0.295
0.216
0.290
0.224
0.303
0.215
0.289
0.223
0.296
0.250
0.309
0.269
0.328
0.281
0.340
0.307
0.376
0.249
0.309
336
0.269
0.322
0.271
0.322
2.568
1.238
0.279
0.330
0.268 0.324
0.281
0.342
0.267
0.326
0.274
0.329
0.311
0.348
0.325
0.366
0.339
0.372
0.325
0.388
0.321
0.351
720
0.351
0.377
0.361
0.380
2.720
1.287
0.359
0.383
0.420
0.422
0.397
0.421
0.352
0.383
0.362
0.385
0.412
0.407
0.421
0.415
0.422
0.419
0.502
0.490
0.408
0.403
Avg
0.248
0.307
0.254
0.310
2.395
1.177
0.256
0.315
0.267
0.322
0.267
0.332
0.249
0.312
0.256
0.317
0.288
0.332
0.304
0.349
0.324
0.368
0.328
0.382
0.291
0.333
Wins
18
23
1
2
0
0
3
5
0
0
0
0
5
1
2
1
0
0
0
0
0
0
0
0
0
0
  • a
    Taken from Wu et al. (2022).