### v3

# xLSTM-Mixer: Multivariate Time Series Forecasting by Mixing via Scalar Memories

Abstract
Time series data is prevalent across numerous fields, necessitating the development of robust and accurate forecasting models. Capturing patterns both within and between temporal and multivariate components is crucial for reliable predictions. We introduce xLSTM-Mixer, a model designed to effectively integrate temporal sequences, joint time-variate information, and multiple perspectives for robust forecasting. Our approach begins with a linear forecast shared across variates, which is then refined by xLSTM blocks. They serve as key elements for modeling the complex dynamics of challenging time series data. xLSTM-Mixer ultimately reconciles two distinct views to produce the final forecast. Our extensive evaluations demonstrate its superior long-term forecasting performance compared to recent state-of-the-art methods while requiring very little memory. A thorough model analysis provides further insights into its key components and confirms its robustness and effectiveness. This work contributes to the resurgence of recurrent models in forecasting by combining them, for the first time, with mixing architectures.
1 Introduction
Figure 1: xLSTM-Mixer provides excellent forecasts with a very low memory footprint while being sufficiently fast. Details are found in Sec. $4.$
Time series are an essential data modality ubiquitous in many critical fields of application, such as medicine (Hosseini et al., 2021), manufacturing (Essien and Giannetti, 2020), logistics (Seyedan and Mafakheri, 2020), traffic management (Lippi et al., 2013), finance (Lin et al., 2012; Divo et al., 2025), and weather modeling (Lam et al., 2023). While significant progress in time series forecasting has been made over the decades, the field is still far from being solved. Further increasing the forecast quality obtained from machine learning models promises a manifold of improvements, such as more accurate medical treatments, increased efficiency in manufacturing and transportation, and higher crop yields.
Historically, recurrent neural networks (RNNs) and their powerful successors were natural choices for deep learning-based time series forecasting (Hochreiter and Schmidhuber, 1997; Cho et al., 2014). Today, large Transformers (Vaswani et al., 2017) are applied extensively to time series tasks, including forecasting. Many improvements to the vanilla architecture have since been proposed, including patching (Nie et al., 2023), decompositions (Zeng et al., 2023), and tokenization inversions (Liu et al., 2023). Newer approaches include pretrained models, which, however, usually cannot capture relationships between multiple variables (Ansari et al., 2024). Yet, some fundamental limitations of Transformers are yet to be lifted. For instance, they are inefficient when applied to long sequences due to the cost of the attention mechanism being quadratic in the number of variates and time steps. As embedded devices and edge computing platforms are gaining importance, the demand for lightweight forecasting models that balance accuracy with minimal memory and computational overhead grows. Therefore, recurrent and state space models (SSMs) (Patro and Agneeswaran, 2025) are experiencing a resurgence of interest in overcoming such limitations. Specifically, Beck et al. (2024) revisited recurrent models by borrowing insights gained from Transformers in many domains, specifically natural language processing. They propose Extended Long Short-Term Memory (xLSTM) models as alternatives to current sequence models.
We propose xLSTM-Mixer [^1] Code available at https://github.com/mauricekraus/xlstm-mixer, a new state-of-the-art method for time series forecasting using recurrent deep learning methods, which strikes a balance by providing strong forecasting accuracy while remaining highly efficient, as Fig. 1 shows. Architecturally, we combine the highly expressive xLSTM architecture with carefully crafted time, variate, and multi-view mixing. These operations regularize the training and limit the number of model parameters by weight-sharing, effectively improving the learning of features necessary for accurate forecasting. xLSTM-Mixer initially computes a channel-independent linear forecast shared over the variates. It is then up-projected to a higher hidden dimension and subsequently refined by an xLSTM stack. It performs multi-view forecasting by producing a forecast from the original and reversed up-projected embedding. The powerful xLSTM cells thereby jointly mix time and variate information to capture complex patterns from the data. Both forecasts are eventually reconciled by a learned linear projection into the final prediction, called view mixing. An overview of our method is shown in Fig. 2.
Figure 2: The xLSTM-Mixer architecture consists of three stages: (1) An initial NLinear forecast assuming channel independence and performing time mixing; (2) subsequent joint mixing, which mixes variate and time information through crucial applications of sLSTM blocks; and (3) view mixing, where the two latent forecast views are reconciled into a coherent final forecast.
Overall, we make the following contributions:
  1. (i)
    We investigate time and variate mixing in the context of recurrent models and propose a joint multistage approach that is highly effective for multivariate time series forecasting. We argue that marching over the variates instead of the temporal axis yields better results if suitably combined with temporal mixing.
  2. (ii)
    We propose xLSTM-Mixer, a state-of-the-art method for time series forecasting, for the first time combining recurrent deep learning with a mixing architecture.
  3. (iii)
    We extensively compare xLSTM-Mixer with existing methods for multivariate long-term time series forecasting and perform in-depth model analyses. The experiments demonstrate that xLSTM-Mixer consistently excels in a wide range of benchmarks.
Outline.   In the upcoming Sec. 2, we introduce preliminaries to allow us to motivate and explain xLSTM-Mixer in Sec. 3. We then comprehensively evaluate its effectiveness and inner workings in Sec. 4. We finally review related work in Sec. 5 and close with a conclusion and outlook in Sec. 6.
2 Background
After introducing the notation used throughout this work, we review xLSTM blocks and discuss whether leveraging channel mixing or their independence is beneficial in time series models.
Notation.   In multivariate time series forecasting, the model is presented with a time series $\bm{X}=\left(\bm{x}_{1},\dots,\bm{x}_{T}\right)\in\mathbb{R}^{V\times T}$ consisting of $T$ time steps with $V$ variates each. Given this context, the forecaster shall predict the future values $\bm{Y}=\left(\bm{x}_{T+1},\dots,\bm{x}_{T+H}\right)\in\mathbb{R}^{V\times H}$ up to a horizon $H$. A variate (sometimes called a channel) can be any scalar measurement, such as the occupancy of a road or the temperature in a power plant. The measurements are assumed to be carried out jointly, such that the $T+H$ time steps reflect a regularly sampled signal. A time series dataset consists of NN such pairs $\left\{\left(\bm{X}^{(i)},\bm{Y}^{(i)}\right)\right\}_{i\in\{1,\dots,N\}}$ divided into train, validation, and test portions.
2.1 Extended Long Short-Term Memory (xLSTM)
Beck et al. (2024) propose xLSTM consisting of two building blocks, namely the sLSTM and mLSTM modules. To harness the full expressivity of xLSTMs within each step and across the computation sequence, we employ a stack of sLSTM blocks without any mLSTM blocks. The latter are less suited for joint mixing due to their independent treatment of the sequence elements, making it impossible to learn any relationships between them directly. See App. B for a deeper discussion. We will continue by recalling how sLSTM cells function. The standard LSTM architecture of Hochreiter and Schmidhuber (1997) involves updating the cell state $\mathbf{c}_{t}$ through a combination of input, forget, and output gates regulating the flow of information across tokens. sLSTM blocks enhance this by incorporating exponential gating and memory mixing (Greff et al., 2017) to handle complex temporal and cross-variate dependencies better. The sLSTM updates the cell $\bm{c}_{t}$ and hidden state $\bm{h}_{t}$ as follows:
$\bm{c}_{t}$ $=\bm{f}_{t}\odot\bm{c}_{t-1}+\bm{i}_{t}\odot\bm{z}_{t}$ cell state (1)
$\bm{n}_{t}$ $=\bm{f}_{t}\cdot\bm{n}_{t-1}+\bm{i}_{t}$ normalizer state (2)
$\bm{h}_{t}$ $=\bm{o}_{t}\odot\bm{c}_{t}\odot\bm{n}^{-1}_{t}$ hidden state (3)
$\bm{z}_{t}$ $=\tanh\bigl(\bm{W}_{z}\bm{x}_{t}+\bm{R}_{z}\bm{h}_{t-1}+\bm{b}_{z}\bigr)$ cell input (4)
$\bm{i}_{t}$ $=\exp\bigl(\bm{\tilde{i}}_{t}-\bm{m}_{t}\bigr)$ $\bm{\tilde{i}}_{t}$ $=\bm{W}_{i}\bm{x}_{t}+\bm{R}_{i}\bm{h}_{t-1}+\bm{b}_{i}$ input gate (5)
$\bm{f}_{t}$ $=\exp\bigl(\bm{\tilde{f}}_{t}+\bm{m}_{t-1}-\bm{m}_{t}\bigr)$ $\bm{\tilde{f}}_{t}$ $=\bm{W}_{f}\bm{x}_{t}+\bm{R}_{f}\bm{h}_{t-1}+\bm{b}_{f}$ forget gate (6)
$\bm{o}_{t}$ $=\sigma\bigl(\bm{W}_{o}\bm{x}_{t}+\bm{R}_{o}\bm{h}_{t-1}+\bm{b}_{o}\bigr)$ output gate (7)
$\bm{m}_{t}$ $=\max\bigl(\bm{\tilde{f}}_{t}+\bm{m}_{t-1},\bm{\tilde{i}}_{t}\bigr)$ stabilizer state (8)
In this setup, the matrices $\bm{W}_{z},\bm{W}_{i},\bm{W}_{f},$ and $\bm{W}_{o}$ are input weights mapping the input token $\bm{x}_{t}$ to the cell input $\bm{z}_{t}$, input gate, forget gate, and output gate, respectively. The states $\bm{n}_{t}$ and $\bm{m}_{t}$ serve as necessary normalization and training stabilization, respectively. As Beck et al. have shown, it is beneficial to restrict the memory mixing performed by the recurrent weight matrices $\bm{R}_{z},\bm{R}_{i},\bm{R}_{f},$ and $\bm{R}_{o}$ to individual heads, inspired by the multi-head setup of Transformers (Zeng et al., 2023), yet more restricted and efficient. In particular, each token gets broken up into groups of features, where the input weights $\bm{W}_{z,i,f,o}$ act across all of them, but the recurrence matrices $\bm{R}_{z,i,f,o}$ are implemented as block-diagonal. This permits specialization of the individual heads to patterns specific to the respective section of the tokens and empirically does not sacrifice expressivity.
2.2 Channel Independence and Mixing for Time Series
Multiple works have investigated whether it is beneficial to learn representations of the time and variate dimensions jointly or separately. Intuitively, because joint mixing is strictly more expressive, one might think it should always be preferred. And, indeed, it is used in many works including Temporal Convolutional Networks (Lea et al., 2016), N-BEATS (Oreshkin et al., 2019), N-HiTS (Challu et al., 2023), and many Transformers (Lim et al., 2021; Wu et al., 2021; Zhou et al., 2022). However, treating slices of the input data independently assumes an invariance to temporal or variate positions and serves as a strong regularization against overfitting, reminiscent of kernels in CNNs. Prominent models implementing some aspects of channel independence in multivariate time series forecasting are PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2023). TiDE (Das et al., 2023), on the other hand, contains a time-step shared feature projection and temporal decoder but treats variates jointly. As Tolstikhin et al. (2021) have shown with MLP-Mixer, interleaving mixing of all channels per token and all tokens per channel does not empirically sacrifice any expressivity and instead improves performance and efficiency. This idea has since been applied to time series too, namely in architectures such as TimeMixer(++) (Wang et al., 2024a, 2025a), TSMixer (Chen et al., 2023c), WPMixer (Murad et al., 2025), and PMformer (Lee et al., 2024), and is, therefore, one key component of our method xLSTM-Mixer.
3 xLSTM-Mixer
We now explain xLSTM-Mixer shown in Fig. 2 in more detail. It carefully integrates several key components: an initial linear forecast with time mixing, joint mixing using powerful sLSTM blocks, and an eventual combination of two views by a final fully connected layer. The transposing steps between the components enable capturing complex temporal and intra-variate patterns while facilitating easy trainability and limiting parameter counts. The sLSTM blocks, in particular, can learn intricate non-linear relationships hidden within the data along both the time and variate dimensions. The architecture is furthermore equipped with normalization layers and skip connections to improve training stability and overall effectiveness.
3.1 Normalization and Initial Linear Forecast
Normalization has become an essential ingredient of modern deep learning architectures (Huang et al., 2023). For time series in particular, reversible instance norm (RevIN) (Kim et al., 2022) is a general recipe for improving forecasting performance, where each time series instance is normalized by its mean and variance and furthermore scaled and offset by a learnable scale $\bm{\gamma}$ and offset $\bm{\beta}$:
$\bm{x}_{t}^{\text{norm}}=\operatorname{RevIN}(\bm{x}_{t})=\bm{\gamma}\odot\left(\frac{\bm{x}_{t}-\mathbb{E}\left[\bm{x}\right]}{\sqrt{\mathrm{Var}\left[\bm{x}\right]}+\bm{\epsilon}}\right)+\bm{\beta}.$
We apply it as part of xLSTM-Mixer, and at the end of the entire pipeline, we invert the RevIN operation to obtain the final prediction. In the case of xLSTM-Mixer, the typical skip connections found in mixer acrchitectures (Tolstikhin et al., 2021; Chen et al., 2023c) are taken up by RevIN, the normalization in the NLinear forecast (Zeng et al., 2023), and the integral skip connections within each sLSTM block.
It has been shown previously that simple linear models equipped with appropriate normalization schemes are, already by themselves, decent long-term forecasters (Zeng et al., 2023; Li et al., 2023). Our observations confirm this finding. Therefore, we first process each variate separately by an NLinear model by computing:
$\bm{x}^{\text{initial}}=\operatorname{NLinear}(\bm{x}^{\text{norm}})=\operatorname{FC}\left(\bm{x}_{1:T}^{\text{norm}}-x_{T}^{\text{norm}}\right)+x_{T}^{\text{norm}}\,$,
where $\operatorname{FC}(\cdot):\mathbb{R}^{T}\to\mathbb{R}^{H}$ denotes a fully-connected linear layer with bias term. Sharing this model across variates limits parameter counts, and the weight-tying serves as a useful regularization. The quality of this initial forecast will be investigated in Sec. 4.1 and 4.5.
3.2 sLSTM Refinement
While the NLinear forecast $\bm{x}^{\text{initial}}\in\mathbb{R}^{V\times H}$ captures the basic patterns between the historic and future time steps, its quality alone is insufficient for today’s challenging datasets. We, therefore, refine it using powerful sLSTM blocks. As a first step, it is crucial to increase the embedding dimension of the data to provide sufficient latent dimensions $D$ for the sLSTM cells as $\bm{x}^{\text{up}}=\operatorname{FC}^{\text{up}}\left(\bm{x}^{\text{initial}}\right)$. This prior up-projection is similar to what is commonly performed in SSMs (Beck et al., 2024). We weight-share $\operatorname{FC}^{\text{up}}:\mathbb{R}^{H}\to\mathbb{R}^{D}$ across variates to perform time-mixing similar to the initial forecast. Note that this step does not maintain the temporal ordering within the embedding token dimensions, as was the case up until this step, and instead embeds it into a higher latent dimension.
The stack of $M$ sLSTM blocks $\mathcal{S}(\cdot)$ transforms $\bm{x}^{\text{up}}\in\mathbb{R}^{V\times D}$ as defined in Eq. 1 to 8. The recurrent model strides over the data in variate order, i.e., where each token represents all time steps from a single variate as in the work of Liu et al. (2023). The sLSTM blocks learn intricate non-linear relationships hidden within the data along both the time and variate dimensions. The mixing of the hidden state is still limited to blocks of consecutive dimensions, aiding efficient learning and inference while allowing for effective cross-variate interaction during the recurrent processing. Striding over variates has the benefit of linear time scaling in the number of variates at a constant number of parameters. It, however, comes at the cost of possibly fixing a suboptimal order of variates. While this is empirically not a significant limitation (see Sec. 4.4), we leave investigations into how to find a suitable ordering for future work. In addition to a large embedding dim, we observed a high number of heads being crucial for effective forecasting.
The sLSTM cells’ first hidden state $\bm{h}_{0}$ must be initialized before each sequence of tokens can be processed. Extending the initial description of these blocks, we propose learning a single initial embedding token $\bm{\eta}\in\mathbb{R}^{D}$ that gets prepended to each encoded time series $\bm{x}^{\text{up}}$. These initial embeddings draw from recent advances in Large Language Models, where learnable "soft prompt" tokens are used to condition models and improve their ability to generate coherent outputs (Lester et al., 2021; Li and Liang, 2021; Chen et al., 2023a, b). Recent research has extended the application of soft prompts to LLM-based time series forecasting (Cao et al., 2023; Sun et al., 2024; Jin et al., 2024), emphasizing their adaptability and effectiveness in improving model performance across modalities. These tokens enable greater flexibility by conditioning its initial memory representation to specific dataset characteristics for dynamically interacting with the time and variate data. Soft prompts can be readily optimized through back-propagation with very little overhead.
3.3 Multi-View Mixing
To further regularize the training of the sLSTM as with the linear projections, we compute forecasts from the original embedding $\bm{x}^{\text{up}}$ as well as the reversed embedding $\bm{\widehat{x}}^{\text{up}}$, where the order of the latent dimensions including the representation of $\bm{\eta}$ is inverted. Learning forecasts $\bm{y}^{\prime},\bm{y}^{\prime\prime}\in\mathbb{R}^{V\times D}$ for both views while sharing weights helps learn better representations. Such multi-task learning settings are known to benefit training (Zhang and Yang, 2022). It can also be viewed as ensembling over different variate orderings with weight sharing. The final forecast is obtained by a linear projection $\operatorname{FC}^{\text{view}}:\mathbb{R}^{D}\times\mathbb{R}^{D}\to\mathbb{R}^{H}$ of the two forecasts, again per-variate. Specifically, we compute $\bm{y}^{\text{norm}}=\operatorname{FC}^{\text{view}}\left(\bm{y}^{\prime},\bm{y}^{\prime\prime}\right)$, where $\bm{y}^{\prime}=\mathcal{S}(\bm{x}^{\text{up}})$ and $\bm{y}^{\prime\prime}=\mathcal{S}(\bm{\widehat{x}}^{\text{up}})$. The final forecast is obtained after de-normalizing the reconciled forecasts as $\bm{y}=\operatorname{RevIN}^{-1}(\bm{y}^{\text{norm}})$.
4 Experiments
We conduct a series of experiments to evaluate the forecasting capabilities of xLSTM-Mixer, aiming to provide comprehensive insights into its performance. Our primary focus is on long-term forecasting, following the works of Das et al. (2023); Chen et al. (2023c); Lin et al. (2024), and Liu et al. (2025). Further tasks are explored in Sec. 4.3. Additionally, we perform an extensive model analysis, including visualizations of the initial embedding tokens, hyperparameter sensitivity, and performance measurement. Finally, an ablation study identifies the contributions of the individual components.
Datasets.   We generally follow the established benchmark procedure of Wu et al. (2021) and Zhou et al. (2021) for best backward and future comparability. The datasets we thus used are provided as an overview in App. D. Training.   We follow standard practice in the forecasting literature by evaluating long-term forecasts using mean squared error (MSE) and mean absolute error (MAE). Based on our experiments, we used MAE as the training loss function since it yielded the best results. The datasets were standardized for consistency across features. In addition, we conducted all experiments three times and reported the averaged values. Further details on hyperparameter selection, metrics, and implementation can be found in App. C. Baseline Models.   We compare xLSTM-Mixer to the recurrent models xLSTMTime (Alharthi and Mahmood, 2024) and LSTM (Hochreiter and Schmidhuber, 1997); mixer models TimeMixer++ (Wang et al., 2025a), TimeMixer (Wang et al., 2024a), and TSMixer (Chen et al., 2023c); MLP-based models CycleNet/MLP (Lin et al., 2024), DLinear (Zeng et al., 2023), and TiDE (Das et al., 2023); the SSMs S-Mamba (Wang et al., 2025b) and Chimera (Behrouz et al., 2024); the Transformers PatchTST (Nie et al., 2023) and iTransformer (Liu et al., 2023); the convolutional architectures ModernTCN (Donghao and Xue, 2023) and TimesNet (Wu et al., 2022a); and the pretrained zero-shot forecasters Timer-XL (Liu et al., 2025) and MoiraiBase (Woo et al., 2024). On choosing lookback lengths $L$.   Some prior works on long-term forecasting fix the lookback window $L$ for the sake of a fair comparison. However, Abdelmalak et al. (2025) and Brigato et al. (2025) argue that fixing especially to 96, which is common in today’s benchmarks, can substantially distort the comparisons. When allowed to tune each model’s input length, most baselines improve and simple linear or MLP backbones close much of the gap to Transformers, whereas at fixed $L=96$ they appear artificially weak. We agree with these findings and therefore provide results for optimal hyperparameters for all baselines and xLSTM-Mixer in our experiments. See also 4(b) in Sec. 4.4 for a sensitivity analysis on $L$.
4.1 Long-Term Time Series Forecasting
Table 1: xLSTM-Mixer is effective in long-term forecasting. The results are averaged from 4 different prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction. The best result for each dataset is highlighted bold red and second-best blue and underlined. Wins for each model out of all 28 settings are shown at the bottom. Full results are provided in App. E.
Models Recurrent Mixer MLP State Space Trans. Conv. Pretrained*
xLSTM-
xLSTMTime
LSTM
TimeMix.++
TimeMix.
TSMixer
CycleNet
DLinear
TiDE
S-Mamba
Chimera
PatchTST
Mod.TCN
Timer-XL
MoiraiBase
Mixer
1997 $\dagger$
Dataset
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Weather
0.219
0.250
0.222 0.255
0.444
0.454
0.226
0.262
0.222
0.262
0.225
0.264
0.223
0.264
0.246
0.300
0.236
0.282
0.251
0.276
0.219
0.258
0.241
0.264
0.224
0.264
0.256
0.294
0.287
0.281
Electricity
0.153
0.245
0.157
0.250
0.559
0.549
0.165
0.253
0.156
0.246
0.160
0.256
0.156
0.251
0.166
0.264
0.159
0.257
0.170
0.265
0.154
0.249
0.159
0.253
0.156
0.253
0.174
0.278
0.187
0.274
Traffic
0.392
0.253
0.391
0.261
1.011
0.541
0.416
0.264
0.387
0.262
0.408
0.284
0.403
0.282
0.434
0.295
0.356
0.261
0.414
0.276
0.403
0.286
0.391
0.264
0.396
0.270
$\ddagger$
$\ddagger$
ETTh1
0.397
0.420
0.408
0.428
1.198
0.821
0.419
0.432
0.411
0.423
0.412
0.428
0.435
0.440
0.423
0.437
0.419
0.430
0.455
0.450
0.405
0.424
0.413
0.434
0.404
0.420
0.404
0.417
0.417
0.419
ETTh2
0.340
0.382
0.346
0.386
3.095
1.352
0.339
0.380
0.316
0.384
0.355
0.401
0.367
0.405
0.431
0.447
0.345
0.394
0.381
0.405
0.318
0.375
0.324
0.381
0.322
0.379
0.347
0.388
0.362
0.382
ETTm1
0.339
0.366
0.347
0.372
1.142
0.782
0.369
0.378
0.348
0.375
0.347
0.375
0.360
0.388
0.357
0.379
0.355
0.378
0.398
0.405
0.345
0.377
0.353
0.382
0.351
0.381
0.373
0.392
0.406
0.385
ETTm2
0.248
0.307
0.254
0.310
2.395
1.177
0.269
0.320
0.256
0.315
0.267
0.322
0.263
0.324
0.267
0.332
0.249
0.312
0.288
0.332
0.250
0.316
0.256
0.317
0.253
0.314
0.273
0.336
0.311
0.337
Wins
11
16
0
2
0
0
0
2
2
2
0
0
1
0
0
0
5
1
0
0
8 3
1
0
2
3
0
1
0
0
    Zero-shot forecasting.
  • $\dagger$
    Taken from Wu et al. (2022a).
  • $\ddagger$
    Traffic/PEMS are often used during pre-training (Liu et al., 2025). Thus, no zero-shot results are available.
Figure 3: xLSTM-Mixer provides convincing forecasts. This figure shows example forecasts on the Weather and ETTm1 datasets for multiple models with lookback windows and forecasting horizons fixed at 96. The first panel illustrates the forecast from xLSTM-Mixer, while the second shows the initial forecast extracted before the up-projection step, highlighting the effectiveness of our added components. Comparisons with further baselines are provided for context.
We present the performance of xLSTM-Mixer compared to prior models in Tab. 1. The full results and standard deviations are provided in App. E. As shown, xLSTM-Mixer consistently delivers highly accurate forecasts across a wide range of datasets. It achieves the best results in 11 out of 28 cases for MSE and 16 out of 28 cases for MAE, demonstrating its superior performance in long-term forecasting. xLSTM-Mixer defines a new state-of-the-art on six out of seven datasets. This shows that xLSTM-Mixer consistently delivers excellent performance in long-term forecasting. A qualitative comparison with several baselines, including the initial forecast extracted before the sLSTM refinement, is shown in Fig. 3. Here, both the lookback window and the forecasting horizon are fixed at 96.
Significance Testing.   In addition to the common practice of evaluating by counting wins across models and datasets, we perform statistical testing following the well-known practice of Demšar (2006). First, a Friedman test ensures the model’s performances follow different distributions ($p<10^{-10}$). We can then perform a Conover post-hoc test while adjusting for the family-wise error rate using Holm’s method. We must restrict this comparison to one metric (MSE) and horizon (96 steps), to avoid the strongly correlated results for different horizons to artificially inflate significance levels. At a significance threshold of $p=0.05$, we find that xLSTM-Mixer is statistically significantly better than all other methods, except for xLSTMTime. Yet, the difference in average rank is still impressive, with 1.5 for xLSTM-Mixer and 4.0 for xLSTMTime. Furthermore, xLSTMTime is statistically inseparable from many other models, namely TimeMixer, TSMixer, ModernTCN, Chimera, PatchTST, and TiDE. See Fig. 5 in App. E for a visual critical difference (CD) diagram.
4.2 Forecasting on the GIFT-Eval Benchmark
To showcase xLSTM-Mixer’s versatility beyond multivariate long-term point forecasting, we evaluate it on GIFT-Eval (Aksu et al., 2024). It is a comprehensive benchmark for general time-series forecasting spanning diverse domains, frequencies, horizons, and both univariate and multivariate regimes. The benchmark emphasizes probabilistic assessment and supports standardized (including zero-shot) evaluation protocols. Following the official setup, we report aggregated Mean Absolute Scaled Error (MASE) and Continuous Ranked Probability Score (CRPS), where lower is better. To extend xLSTM-Mixer to probabilistic forecasting capabilities, we equip it with a head predicting quantiles, yielding full predictive distributions rather than point estimates.
Results are summarized in Tab. 2 and detailed in App. F. By CRPS, xLSTM-Mixer is the top purely supervised model (i.e., without using the GIFT-Eval pretraining corpus) and ranks 2nd overall. This demonstrates that although xLSTM-Mixer is designed for multivariate forecasting, it remains highly competitive among methods across univariate and multivariate settings and for both short and long horizons in the probabilistic regime.
Table 2: xLSTM-Mixer excels on GIFT-Eval. The table reports the top 10 across all categories.
Model MASE $\downarrow$ CRPS $\downarrow$ Rank (CRPS) $\downarrow$
TiRex $0.724$ $0.498$ 11
xLSTM-Mixer (ours) $0.780$ $0.510$ 22
TEMPO_ensemble $0.862$ $0.514$ 33
Toto_Open_Base_1.0 $0.750$ $0.517$ 44
TabPFN-TS $0.771$ $0.544$ 55
YingLong_300m $0.798$ $0.548$ 66
timesfm_2_0_500m $0.758$ $0.550$ 77
YingLong_110m $0.809$ $0.557$ 88
sundial_base_128m $0.750$ $0.559$ 99
YingLong_50m $0.822$ $0.567$ $10$
Significance Testing.   As in the previous Sec. 4.1, we rigorously test for the significance of the improvement of xLSTM-Mixer over the existing methods. After the initial Firedman test on the MSE, we see that xLSTM-Mixer is significantly better than all other 38 methods, except for sundial_base_128m, TiRex, and TTM-R2-Finetuned. See Fig. 6 in App. F for a CD diagram.
4.3 Outlook on Classification
To illustrate performance beyond forecasting, we evaluate xLSTM-Mixer on time-series classification by using it as an embedding model. For this, we replace the final projection layer with a single classification head. xLSTM-Mixer achieves very strong performance on standard time-series classification benchmarks (cf. App. J). These findings suggest that approaches reconciling recurrent and mixing architectures, such as xLSTM-Mixer, are highly flexible yet powerful.
4.4 Model Analysis
Initial Token Embeddings.   We qualitatively inspect decodings of the initial embedding tokens $\bm{\eta}$ on multiple datasets to further understand and interpret the initializations learned by xLSTM-Mixer. $\bm{\eta}$ are decoded to a forecast $\bm{y}$ by transforming them through the sLSTM stack $\mathcal{S}$ and applying multi-view mixing. The resulting output of $\operatorname{FC}^{\text{view}}$ can then be interpreted as the conditioning time series used to initialize the sLSTM blocks. Fig. 7 in Appendix App. G shows the dataset-specific patterns the initial embedding tokens have learned for various horizons $H$. Increasing prediction horizons eventually reveal underlying seasonal patterns and respective dataset dynamics.
Model Efficiency.   To assess the computational resources required for using xLSTM-Mixer, we measured the average wall-clock time and peak graphics card memory required to perform a training step. Fig. 1 shows how this changes over multiple lookback lengths $T$ and two datasets at a forecast horizon of $H=336$. Compared to baselines, xLSTM-Mixer scales extremely favorably in $T$, only exhibiting a negligible increase in time and memory requirements. While computations take slightly longer for larger lookback sizes, the increase is much smaller than for Transformer-based models. One advantage of TimeMixer was its efficiency over Transformers, upon which xLSTM-Mixer now significantly improves by requiring one to two orders of magnitude less memory.
Sensitivity to the xLSTM Hidden Dimension.   In 4(a), we visualize the performance of xLSTM-Mixer on the Electricity dataset with increasing sLSTM embedding hidden dimension $D$ realized by $\operatorname{FC}^{\text{up}}$. The results indicate that a larger $D$ enables xLSTM-Mixer to better capture the complexity of the series over extended horizons, leading to improved forecasting accuracy.
(a) Increasing the hidden dimension $D$ becomes increasingly beneficial at longer prediction lengths $H$. Shown is the MSE for varying $D$ on the Electricity dataset.
(b) Increasing the lookback window $T$ increases forecasting performance, with xLSTM-Mixer virtually always providing the best results. Shows the MSE (↓) on the ETTm1 dataset.
Figure 4: Impact of model parameters on forecasting performance.
Robustness to the Lookback Length.   4(b) illustrates the performance of xLSTM-Mixer across varying lookback lengths $T$ and prediction horizons $H$. Note that we had to rerun some experiments for TimeMixer at $T=720$ with varying seeds since many training runs diverged. We observe that xLSTM-Mixer can effectively utilize longer lookback windows than the baselines, especially when compared to transformer-based models. This advantage stems from xLSTM-Mixer’s avoidance of self-attention, allowing it to handle extended lookback lengths efficiently. On short prediction lengths with $T\in\{96,192\}$, information of more than 768 time steps in the past becomes redundant to inform the comparatively short forecast, causing models to deteriorate slightly. On longer horizons, increasingly farther lookbacks become useful for forecasting. Additionally, xLSTM-Mixer demonstrates stable and consistent performance with low variance across scales.
xLSTM-Mixer Captures Cross-Variate Patterns.   To determine if our chosen ordering of variates is viable, we conduct additional experiments investigating the impact of different variate permutations in Tab. 12 in App. H. While we observe that the performance does depend to a certain extent on the ordering of variates, this does not pose a significant limitation in practice since the standard ordering provided with each dataset already yields strong forecasts. Moreover, our attribution analysis (cf. Fig. 8) confirms that xLSTM-Mixer effectively captures cross-variate interactions.
Ensembling over More than Two Views.   Following the perspective that multi-view mixing is an ensemble over variate orderings, we extend our method to accommodate additional randomly permuted views beyond the original and reverse order. To this end, $\operatorname{FC}^{\text{view}}$ is extended accordingly. The full results in App. I show that this does not improve the modeling accuracy of xLSTM-Mixer.
4.5 Ablation Studies
Table 3: Each component of xLSTM-Mixer is essential for its overall strong performance. The notation follows Tab. 1. Results are averages over three seeds.
#1 (full)
#2
#3
#4
#5
#6
#7
#8
#9
#10
#11
#12
#13
#14
Time Mixing
DLinear
xLSTM type
sLSTM
mLSTM
LSTM
GRU
sLSTM
sLSTM
sLSTM
sLSTM
sLSTM
None
sLSTM
sLSTM
sLSTM
sLSTM
Recurr. order
Variates
Variates
Variates
Variates
Time
Variates
Variates
Variates
Variates
None
Variates
Variates
Variates
Variates
Init. Token
View Mixing
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Weather
96
0.143
0.184
0.148
0.192
0.148
0.190
0.147
0.193
0.148
0.194
0.145
0.187
0.145
0.186
0.144
0.185
0.144
0.186
0.173
0.223
0.149
0.193
0.151
0.195
0.149
0.192
0.152
0.195
192
0.186
0.226
0.193
0.235
0.191
0.232
0.188
0.233
0.196
0.239
0.188
0.229
0.188
0.228
0.186
0.226
0.188
0.228
0.219
0.257
0.192
0.233
0.192
0.234
0.191
0.234
0.193
0.236
336
0.237
0.266
0.241
0.272
0.241
0.271
0.238
0.272
0.252
0.281
0.237
0.267
0.239
0.267
0.241
0.270
0.242
0.270
0.261
0.288
0.240
0.271
0.242
0.273
0.242
0.273
0.244
0.274
720
0.310
0.324
0.313
0.325
0.327
0.344
0.345
0.361
0.315
0.328
0.312
0.325
0.310
0.324
0.309
0.323
0.309
0.323
0.320
0.334
0.320
0.329
0.319
0.329
0.322
0.330
0.319
0.328
ETTm1
96
0.275
0.328
0.285
0.339
0.317
0.367
0.312
0.361
0.298
0.348
0.274
0.329
0.277
0.329
0.278
0.331
0.279
0.333
0.295
0.338
0.282
0.339
0.285
0.341
0.281
0.337
0.284
0.339
192
0.319
0.354
0.329
0.365
0.354
0.388
0.349
0.382
0.337
0.369
0.319
0.356
0.321
0.354
0.321
0.356
0.322
0.358
0.329
0.357
0.329
0.364
0.330
0.365
0.337
0.367
0.335
0.366
336
0.353
0.374
0.363
0.384
0.383
0.403
0.375
0.395
0.368
0.388
0.351
0.376
0.354
0.375
0.355
0.377
0.357
0.379
0.359
0.376
0.367
0.385
0.367
0.385
0.366
0.384
0.366
0.385
720
0.409
0.407
0.417
0.414
0.440
0.432
0.447
0.441
0.420
0.416
0.409
0.408
0.411
0.408
0.413
0.411
0.414
0.411
0.412
0.407
0.422
0.412
0.422
0.413
0.417
0.410
0.418
0.411
To assess the contributions of each component in xLSTM-Mixer to its strong forecast performance, we conducted ablation studies of thirteen different model configurations with the results listed in Tab. 3. Each configuration represents a different combination of the four key components: mixing time with NLinear/DLinear, using sLSTM/mLSTM/LSTM/GRU for joint mixing, striding over the variate or time dimension, learning initial embedding tokens $\bm{\eta}$, and multi-view mixing.
The full version of xLSTM-Mixer (#1), which integrates all components, achieves the best performance overall. Ablating components of xLSTM-Mixer each causes both error metrics to increase, entailing that they contribute positively to the overall architecture. Specifically, omitting xLSTM (only having LSTM, #3) raises the MAE by 6.2%/the MSE by 7.0%, variate recurrence (#5) by 4.3%/4.7%, learning initial token embedding (#7) by 0.4%/0.7%, view mixing (#8) by 0.6%/0.7%, and time mixing (#11) by 2.7%/3.1%. For example, removing the time mixing (#11) increases the MAE by 3.4% on ETTm1 at length 96 or 3.1% at length 192, highlighting its critical role in capturing cross-time dependencies. When we now omit everything except for time mixing on Weather at 192, we suffer a 13.7% performance decrease. Additionally, substituting sLSTM blocks with mLSTM, LSTM, or GRU (#2–#4) consistently degrades performance, particularly for LSTM and GRU at longer horizons. We attribute this to the sLSTM’s inherent structure, which provides stronger mixing capabilities and mitigates the degradation commonly observed over extended prediction windows. However, we also observe that some configurations of xLSTM-Mixer, which exclude specific components, remain competitive. For instance, #7, which excludes the initial embedding token, still performs very well. This suggests that while it contributes positively to the overall performance, the model can sometimes still achieve competitive results without it. Similarly, depending on the dataset and target metric, initial forecasting with DLinear instead of NLinear is a viable option, too (#6).
The ablation study confirms that all components of xLSTM-Mixer contribute to its effectiveness, with the full configuration yielding the best results. Furthermore, we identified the sLSTM blocks and time-mixing components as critical for ensuring high accuracy across datasets and prediction lengths.
5 Related Work
Time Series Forecasting.   A long line of machine learning research led from early statistical methods like ARIMA (Sims, 1980; Box et al., 2015) to contemporary models based on deep learning, where six architectural families take center stage: based on recurrence, state spaces, convolutions, Multilayer Perceptrons (MLPs), mixing, and Transformers. While all of them are used by practitioners today, the research focus for long-term time series forecasting is gradually shifting over time. Initially, the naturally sequential recurrent models such as Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997) and Gated Recurrent Units (GRUs) (Cho et al., 2014) were used for time series analysis. Their main benefits are the high inference efficiency and arbitrary input and output lengths due to their autoregressive nature. While their effectiveness has historically been constrained by a limited ability to capture long-range dependencies, active research remains to alleviate these limitations (Salinas et al., 2020), including the xLSTM architecture presented in Sec. 2 (Beck et al., 2024; Alharthi and Mahmood, 2024) and SutraNets (Bergsma et al., 2023). Closely related are state space models (SSMs) such as Mamba (Gu and Dao, 2024; Wang et al., 2025b) or Chimera (Behrouz et al., 2024), which permit parallel inference for improved efficiency. Similarly efficient, yet more restricted in their output length, are the location-invariant CNNs (Li et al., 2022; Lara-Benítez et al., 2021), such as TCN (Lea et al., 2016), TimesNet (Wu et al., 2022a), and MICN (Wang et al., 2022). Recently, some MLP-based architectures have also shown good success, including the simplistic DLinear and NLinear models (Zeng et al., 2023), the encoder-decoder architecture of TiDE (Das et al., 2023), and the older hierarchical N-BEATS (Oreshkin et al., 2019) and N-HiTS (Challu et al., 2023) models. They are closely related to other mixing architectures such as TimeMixer(++) (Wang et al., 2024a, 2025a) and TSMixer (Chen et al., 2023c). A lot of accurate models with significant compute costs have been proposed based on Transformers (Vaswani et al., 2017), such as Autoformer (Wu et al., 2021), TFT (Lim et al., 2021), FEDFormer (Zhou et al., 2022), PatchTST (Nie et al., 2023), and iTransformer (Liu et al., 2023). Finally, the most recent development are pretrained models fitted on multiple time series datasets (Kraus et al., 2024). Examples include Chronos (Ansari et al., 2024), Moirai (Woo et al., 2024), and Timer-XL (Liu et al., 2025). xLSTM-Mixer combines two model families, namely the highly expressive mixing models with efficient recurrence, to benefit from the strengths of both.
xLSTM Models for Time Series.   Some initial experiments of applying xLSTMs (Beck et al., 2024) to time series were already performed by Alharthi and Mahmood (2024) with their proposed xLSTMTime model. While it showed promising forecasting performance, these initial soundings did not surpass stronger models at that time, such as TimeMixer (Wang et al., 2024a), on multivariate benchmarks. Furthermore, despite our best efforts, the experimental findings unfortunately could not be replicated with either the official implementation or the scarce details in the paper. We ensure that our method xLSTM-Mixer is well-suited as a foundation for further research by providing extensive model analysis, an ablation study with 13 variations, and ensuring that results are readily reproducible. Our methodology draws from xLSTMTime yet improves on it by several key components. Most importantly, our novel multi-view mixing consistently enhances forecasting performance. Furthermore, we find the trend-seasonality decomposition redundant and a simple NLinear normalization scheme (Zeng et al., 2023) to suffice. Concurrently, Kong et al. (2024) also investigate using xLSTMs for time series forecasting, arriving at similar conclusions. Poonia et al. (2025) successfully employ xLSTMs to detect Granger Causality in time series.
6 Conclusion
In this work, we introduced xLSTM-Mixer, a method that combines a linear forecast with refinement through xLSTM blocks. Our architecture integrates time, joint, and view mixing to capture complex dependencies. In long-term forecasting, xLSTM-Mixer consistently achieves state-of-the-art performance, outperforming a large set of previous methods in 27 out of 56 cases. We also evaluate xLSTM-Mixer in heterogeneous probabilistic settings on GIFT-Eval, where it delivers strong, competitive performance. Beyond forecasting, xLSTM-Mixer performs strongly on time-series classification when used as an embedding model. Moreover, xLSTM-Mixer attains these results at a low memory footprint. Our model analysis provided insights into the contribution of each component and demonstrated robustness to hyperparameter variations.
Limitations.   While xLSTM‑Mixer achieves strong accuracy with limited compute, certain assumptions need to be met for it to be applicable. Specifically, it assumes that all variables are sampled on a uniform time grid, meaning that irregular or missing timestamps must still be handled in pre‑processing. Furthermore, treating variates as the sequential axis ties runtime and memory directly to the number of channels, which can become a bottleneck for extremely high‑dimensional multivariate time‑series. Moreover, the simultaneous mixing of multiple views blends temporal and cross‑channel information in ways that make detailed attributions difficult.
Future Work.   Addressing these aspects through adaptive variate grouping, continuous‑time embeddings, and lightweight explanation modules are potential paths forward beyond the current work. Finally, extending xLSTM-Mixer to tasks such as imputation or anomaly detection offers promising future directions.
Acknowledgements
This work received funding from the EU project EXPLAIN, under the German Federal Ministry of Research, Technology and Space (BMFTR) (grant 16IS22030D). Furthermore, it was funded by the ACATIS Investment KVG mbH project “Temporal Machine Learning for Long-Term Value Investing” and the BMFTR KompAKI project within the “The Future of Value Creation – Research on Production, Services and Work” program (funding number 02L19C150), managed by the Project Management Agency Karlsruhe (PTKA). The author of Eindhoven University of Technology received support from their Department of Mathematics and Computer Science and the Eindhoven Artificial Intelligence Systems Institute. Furthermore, this work benefited from the HMWK project “The Third Wave of Artificial Intelligence - 3AI”, the BMFTR project “XEI” (FKZ 16IS24079B), and from early stages of the Clusters of Excellence “Reasonable AI” (EXC-3057) and “The Adaptive Mind” (EXC-3066) funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy; funding will begin in 2026. The authors are responsible for the content of this publication.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We explicitly list the contributions at the end of Sec. 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We explicitly discuss limitations and suggest paths forward at the end of Sec. 6. We acknowledge that any empirical results are by their very nature limited to the settings in which they were obtained, and thus strive to accurately describe them for best reproducibility (see also Question 4).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include mathematical theorems or proofs thereof.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All implementation details, including dataset descriptions in, metric calculations, and experiment configurations, are provided in Sec. 4 and App. D & C.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We make sure to exclusively use openly available software and datasets and provide the source code for full reproducibility at https://github.com/mauricekraus/xlstm-mixer.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All implementation details, including dataset descriptions in, metric calculations, and experiment configurations, are provided in Sec. 4 and App. D & C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All main results are accompanied by standard deviations. See also Tab. 7 in App. E.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: This is provided in App. C.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We made sure to comply with the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We dedicate App. A to this.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Please see Question 10. Specifically, we do not provide any trained models or similar high-risk artifacts.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We carefully cite all immediately relevant scholarly works and provide URLs to any other resources in Sec. 4 and App. C.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., $C$-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide any such information beyond the paper at https://github.com/mauricekraus/xlstm-mixer.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We did not perform any such experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Please see Question 14.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We did not employ LLMs in any part of xLSTM-Mixer.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Societal Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Our research advances machine learning by enhancing the capabilities of long-term forecasting in time series models, significantly improving both accuracy and efficiency. By developing xLSTM-Mixer, we introduce a robust framework that can be applied across various industries, including finance, healthcare, energy, and logistics. The improved forecasting accuracy enables better decision-making in critical areas, such as optimizing resource allocation, predicting market trends, and managing risk.
However, we also recognize the potential risks associated with the misuse of these advanced models. Time series forecasting models could be leveraged for malicious purposes, especially when applied at scale. For example, in the financial sector, adversarial agents might manipulate forecasts to create market instability. In political or social contexts, these models could be exploited to predict and influence public opinion or destabilize economies. Additionally, the application of these models in sensitive domains like healthcare and security may lead to unintended consequences if not carefully regulated and ethically deployed.
Therefore, it is essential that the use of xLSTM-Mixer, like all machine learning technologies, is guided by responsible practices and ethical considerations. We encourage stakeholders to adopt rigorous evaluation processes to ensure fairness, transparency, and accountability in its deployment, and to remain vigilant to the broader societal implications of time series forecasting technologies.
Appendix B Rationale for Employing sLSTM over mLSTM
Beck et al. [2024] present a comprehensive case for xLSTM blocks over earlier recurrent models. Here, we explain why, within those, we choose scalar-memory sLSTM blocks as our mixing primitive rather than mLSTM. The primary motivation is the state tracking via memory mixing. sLSTM preserves hidden-to-hidden recurrences in its gates and cell update, enabling conditional, history-aware updates. mLSTM removes these paths to allow full parallelization, which limits this state tracking capability.
Mechanisms of sLSTM and mLSTM.   The sLSTM block maintains the LSTM-style memory mixing, augmented with exponential gates and a stabilization term:
$c_{t}=f_{t}c_{t-1}+i_{t}z_{t},\qquad\qquad h_{t}=o_{t}g(c_{t}).$ (9)
Here, $f_{t},i_{t},o_{t}$ each depend on $h_{t-1}$ through recurrent matrices (Beck et al. [2024], Greff et al. [2017]), allowing the block to blend past and present conditionally. By contrast, mLSTM writes key-value pairs into a matrix memory:
$C_{t}=f_{t}C_{t-1}+i_{t}v_{t}k_{t}^{\top},$ (10)
with gates that depend only on the current input (no hidden-to-hidden paths) [Beck et al., 2024, Sec. 2.3]. This design eliminates memory mixing. Note that we use the scalar formulation for simplicity, whereas Sec. 2.1 directly presented the full vector-valued xLSTMs.
Importance of Latent Phases.   In long-horizon forecasting, tracking latent phases (e.g., heating $\leftrightarrow$ cooling; up $\leftrightarrow$ down) is critical. With sLSTM, gates can be trained such that, upon detection of a switch cue, the cell overwrites ($i_{t}\approx 1,f_{t}\approx 0$); otherwise, it retains ($i_{t}\approx 0,f_{t}\approx 1$). Because the gates read $h_{t-1}$, the decision is conditional on the current internal state. In mLSTM, gates cannot use $h_{t-1}$, meaning updates become fixed linear functionals of the input history. Empirically, architectures lacking memory mixing (mLSTM, Mamba, Transformers) fail simple state-tracking tasks such as parity, as shown by Merrill et al. [2024] and Beck et al. [2024].
Design Consequences.   For long-horizon forecasting, state-tracking capacity outweighs raw matrix-memory capacity, precluding the direct learning of inter-token relations during joint mixing. We, thus, exclusively employ sLSTM in xLSTM-Mixer. In our ablations, replacing sLSTM with mLSTM, GRU, or LSTM consistently harms long-horizon accuracy, particularly at 336/720 steps (cf. Tab. 3). Prior and concurrent work supports this finding: xLSTMTime uses sLSTM on smaller forecasting datasets where precise phase tracking dominates and opts for mLSTM only when raw storage capacity is paramount [Alharthi and Mahmood, 2024]. TiRex likewise attributes gains to retaining state tracking [Auer et al., 2025].
Appendix C Implementation Details
Experimental Details.
Our codebase is implemented in Python 3.11, leveraging PyTorch version 2. [Paszke et al., 2019] in combination with Lightning version 2.4222https://lightning.ai/pytorch-lightning for model training and optimization. We used the custom CUDA implementation [^3] https://github.com/NX-AI/xlstm for sLSTM, which relies on NVIDIA Compute Capability $\geq$ 8.0. Thus, our experiments were conducted on a single NVIDIA A100 80GB GPU. The majority of our baseline implementations, along with data loading and preprocessing steps, are adapted from the Time-Series-Library [^4] https://github.com/thuml/Time-Series-Library of Wang et al. [2024b]. For GIFT-Eval [Aksu et al., 2024], we integrated xLSTM-Mixer into the official evaluation harness via a GluonTS-style [Alexandrov et al., 2020] estimator. For xLSTMTime, we used code based on the official repository [^5] https://github.com/muslehal/xLSTMTime. We employ Captum [^6] https://captum.ai  [Kokhlikyan et al., 2020] to compute the SHAP values used in model analysis. We used scikit-posthocs [Terpilowski, 2019] for significance analyses.
Training and Hyperparameters.
We optimized xLSTM-Mixer in 32 bits for up to 60 epochs with a cosine-annealing scheduler with the Adam optimizer [Kingma and Ba, 2015], using $\beta_{1}=0.9$ and $\beta_{2}=0.999$ and no weight decay. Hyperparameter (HP) tuning was conducted using Optuna [Akiba et al., 2019] with the choices provided in Tab. 4. We optimized for the L1 forecast error, also known as the Mean Absolute Error (MAE). To further stabilize the training process, gradient clipping with a maximum norm of $1.0$ was applied. All experiments were run with the three different random seeds $\{2021,2022,2023\}$. For most models, the initial publications already provide HPs appropriate for the well-known datasets, where we thus directly adopt these results. For xLSTMTime, this was not the case, and we were not able to fully reproduce the results in the paper despite our best efforts (cf. Sec. 5). We still present the better results from Alharthi and Mahmood [2024] so as not to erroneously underestimate the method. For the datasets ETTh2, ETTm1, and ETTm2 without such results, we ran individual HP searches analogously to our method to ensure a fair comparison.
Table 4: Hyperparameters and their choices.
Hyperparameter Choices
Batch size {16, 32, 64, 128, 256, 512}
Initial learning rate $1\cdot 10^{-2}$, $3\cdot 10^{-3}$, $1\cdot 10^{-3}$, $5\cdot 10^{-4}$, $2\cdot 10^{-4}$, $1\cdot 10^{-4}$}
Scheduler warmup steps {5, 10, 15}
Lookback length {96, 256, 512, 768, 1024, 2048}
Embedding dimension $D$ {32, 64, 128, 256, 512, 768, 1024}
sLSTM conv. kernel width {disabled, 2, 4}
sLSTM dropout rate {0.1, 0.25}
# sLSTM blocks $M$ {1, 2, 3, 4}
# sLSTM heads {4, 8, 16, 32}
Forecasting Metrics.
We follow common practice in the literature [Wu et al., 2021, Wang et al., 2024a] for maximum comparability and, therefore, evaluate deterministic long-term forecasting of all models on the mean absolute error (MAE), mean squared error (MSE). For probabilistic evaluation (as in GIFT-Eval [Aksu et al., 2024]), we use the Mean Absolute Scaled Error (MASE) and the Continuous Ranked Probability Score (CRPS). All metrics are computed per variate and averaged over all variates. Formally,
[…]
The CRPS for predictive CDF $\hat{F}$ and outcome $y$ is
$\operatorname{CRPS}(\hat{F},y)$ $=\int_{-\infty}^{\infty}\!\big(\hat{F}(z)-\mathbbm{1}\{y\leq z\}\big)^{2}\,\mathrm{d}z\;=\;\mathbb{E}\lvert X-y\rvert\;-\;\tfrac{1}{2}\mathbb{E}\lvert X-X^{\prime}\rvert$,
where $\bm{y}$ are the targets, $\hat{\bm{y}}$ the predictions; $H$ is the horizon length, $T$ the in-sample length, $m$ a seasonal period, $d$ the MASE scaling constant (computed per variate), $\hat{F}$ the predictive CDF, and $X,X^{\prime}\!\sim\!\hat{F}$ i.i.d. When forecasts are provided as quantiles $\{\hat{q}_{t,\tau_{k}}\}_{k=1}^{K}$, we use the standard discrete approximation to CRPS via the pinball loss:
$\widehat{\operatorname{CRPS}}_{t}$ $=\frac{2}{K}\sum_{k=1}^{K}\rho_{\tau_{k}}\!\big(y_{t}-\hat{q}_{t,\tau_{k}}\big),\qquad\rho_{\tau}(u)=(\tau-\mathbb{1}\{u<0\})\,u,$
$\overline{\operatorname{CRPS}}$ $=\frac{1}{H}\sum_{t=1}^{H}\widehat{\operatorname{CRPS}}_{t}.$
Appendix D Benchmark Datasets
Tab. 5 provides an overview of the datasets we used to compare xLSTM-Mixer with other time series forecasting models. The last column shows the range of Hurst exponents [Hurst, 1951] over the variates measuring long-term patterns. The larger the values are over 0.5, the more long-term patterns are in the time series.
Table 5: The long-term forecasting benchmark datasets and their key properties.
Dataset Source Domain Horizons Sampling #Variates Hurst exp.
Weather Zhou et al. [2021] Weather 96–720 10 min 21 0.333–1.000
Electricity Zhou et al. [2021] Power Usage 96–720 1 hour 321 0.555–1.000
Traffic Wu et al. [2021] Traffic Load 96–720 1 hour 862 0.162–1.000
ETT Zhou et al. [2021] Power Production 96–720 15&60 min 7 0.906–1.000
Appendix E Full Results for Long-Term Forecasting
Table 6: Full long-term forecasting results for Tab. 1. Avg is averaged from all four prediction lengths {96, 192, 336, 720}. A lower MSE or MAE indicates a better prediction. The best result for each dataset is highlighted bold red and second-best blue and underlined. Wins for each model out of all 28 settings are shown at the bottom.
Models Recurrent Mixer MLP State Space Transformer Convolutional Pretrained*
xLSTM-
xLSTMTime
LSTM
TimeMix.++
TimeMix.
TSMixer
CycleNet
DLinear
TiDE
S-Mamba
Chimera
PatchTST
iTransf.
Mod.TCN
TimesNet
Timer-XL
MoiraiBase
Mixer
1997 $\dagger$
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Weather
96
0.143
0.184
0.144 0.187
0.369
0.406
0.155
0.205
0.147
0.197
0.145
0.198
0.148
0.200
0.176
0.237
0.166
0.222
0.165
0.210
0.146
0.206
0.149
0.198
0.174
0.214
0.149
0.200
0.172
0.220
0.171
0.225
0.220
0.217
192
0.186
0.226
0.192
0.236
0.416
0.435
0.201
0.245
0.189
0.239
0.191
0.242
0.190
0.240
0.220
0.282
0.209
0.263
0.214
0.252
0.189
0.239
0.194
0.241
0.221
0.254
0.196
0.245
0.219
0.261
0.221
0.271
0.271
0.259
336
0.236
0.266 0.237
0.272
0.455
0.454
0.237
0.265
0.241
0.280
0.242
0.280
0.242
0.283
0.265
0.319
0.254
0.301
0.274
0.297
0.244
0.281
0.306
0.282
0.278
0.296
0.238
0.277
0.280
0.306
0.274
0.311
0.286
0.297
720
0.310 0.323
0.313
0.326
0.535
0.520
0.312
0.334
0.310
0.330
0.320
0.336
0.312
0.333
0.323
0.362
0.313
0.340
0.350
0.345
0.297
0.309
0.314
0.334
0.358
0.347
0.314
0.334
0.365
0.359
0.356
0.370
0.373
0.354
Avg
0.219
0.250
0.222 0.255
0.444
0.454
0.226
0.262
0.222
0.262
0.225
0.264
0.223
0.264
0.246
0.300
0.236
0.282
0.251
0.276
0.219
0.258
0.241
0.264
0.258
0.278
0.224
0.264
0.259
0.287
0.256
0.294
0.287
0.281
Electricity
96
0.126
0.218
0.128 0.221
0.375
0.437
0.135
0.222
0.129
0.224
0.131
0.229
0.126
0.221
0.140
0.237
0.132
0.229
0.139
0.235
0.132
0.234
0.129
0.222
0.148
0.240
0.129
0.226
0.168
0.272
0.141
0.237
0.160
0.250
192
0.144
0.235
0.150
0.243
0.442
0.473
0.147
0.235
0.140
0.220
0.151
0.246
0.144
0.237
0.153
0.249
0.147
0.243
0.159
0.255
0.144
0.223
0.147
0.240
0.162
0.253
0.143
0.239
0.184
0.289
0.159
0.254
0.175
0.263
336
0.157 0.250
0.166
0.259
0.439
0.473
0.164
0.245
0.161
0.255
0.161
0.261
0.159
0.255
0.169
0.267
0.161
0.261
0.176
0.272
0.156
0.259
0.163
0.259
0.178
0.269
0.161
0.259
0.198
0.300
0.177
0.272
0.187
0.277
720
0.183
0.276
0.185
0.276
0.980
0.814
0.212
0.310
0.194
0.287
0.197
0.293
0.196
0.290
0.203
0.301
0.196
0.294
0.204
0.298
0.184 0.280
0.197
0.290
0.225
0.317
0.191
0.286
0.220
0.320
0.219
0.308
0.228
0.309
Avg
0.153
0.245
0.157
0.250
0.559
0.549
0.165
0.253
0.156
0.246
0.160
0.256
0.156
0.251
0.166
0.264
0.159
0.257
0.170
0.265
0.154
0.249
0.159
0.253
0.178
0.270
0.156
0.253
0.192
0.295
0.174
0.278
0.187
0.274
Traffic
96
0.357
0.236
0.358
0.242
0.843
0.453
0.392
0.253
0.360
0.249
0.376
0.264
0.374
0.268
0.410
0.282
0.336
0.253
0.382
0.261
0.366
0.248
0.360
0.249
0.395
0.268
0.368
0.253
0.593
0.321
$\ddagger$
$\ddagger$
192
0.377
0.241
0.378
0.253
0.847
0.453
0.402
0.258
0.375 0.250
0.397
0.277
0.390
0.275
0.423
0.287
0.346
0.257
0.396
0.267
0.394
0.292
0.379
0.256
0.417
0.276
0.379
0.261
0.617
0.336
336
0.394
0.250
0.392
0.261
0.853
0.455
0.428
0.263
0.385
0.270
0.413
0.290
0.405
0.282
0.436
0.296
0.355
0.260
0.417
0.276
0.409
0.311
0.392
0.264
0.433
0.283
0.397
0.270
0.629
0.336
720
0.439
0.283
0.434
0.287
1.500
0.805
0.441
0.282
0.430 0.281
0.444
0.306
0.441
0.302
0.466
0.315
0.386
0.273
0.460
0.300
0.443
0.294
0.432
0.286
0.467
0.302
0.440
0.296
0.640
0.350
Avg
0.392
0.253
0.391
0.261
1.011
0.541
0.416
0.264
0.387
0.262
0.408
0.284
0.403
0.282
0.434
0.295
0.356
0.261
0.414
0.276
0.403
0.286
0.391
0.264
0.428
0.282
0.396
0.270
0.620
0.336
ETTh1
96
0.359
0.386
0.368
0.395
1.044
0.773
0.361
0.403
0.361
0.390
0.361
0.392
0.382
0.403
0.375
0.399
0.375
0.398
0.386
0.405
0.362
0.391
0.370
0.400
0.386
0.405
0.368
0.394
0.384
0.402
0.369
0.391
0.376
0.392
192
0.402
0.417
0.401
0.416
1.217
0.832
0.416
0.441
0.409
0.414
0.404
0.418
0.421
0.426
0.405
0.416
0.412
0.422
0.443
0.437
0.398
0.415
0.413
0.429
0.441
0.436
0.405
0.413
0.436
0.429
0.405
0.413
0.412
0.413
336
0.408
0.429
0.422
0.437
1.259
0.841
0.430
0.434
0.430
0.429
0.420
0.431
0.449
0.444
0.439
0.443
0.435
0.433
0.489
0.468
0.402 0.416
0.422
0.440
0.487
0.458
0.391
0.412
0.491
0.469
0.418
0.423
0.433
0.428
720
0.419
0.448
0.441
0.465
1.271
0.838
0.467
0.451
0.445
0.460
0.463
0.472
0.486
0.487
0.472
0.490
0.454
0.465
0.502
0.489
0.458
0.477
0.447
0.468
0.503
0.491
0.450
0.461
0.521
0.500
0.423
0.441
0.447
0.444
Avg
0.397
0.420
0.408
0.428
1.198
0.821
0.419
0.432
0.411
0.423
0.412
0.428
0.435
0.440
0.423
0.437
0.419
0.430
0.455
0.450
0.405
0.424
0.413
0.434
0.454
0.448
0.404
0.420
0.458
0.450
0.404
0.417
0.417
0.419
ETTh2
96
0.267
0.329
0.273
0.333
2.522
1.278
0.276
0.328
0.271
0.330
0.274
0.341
0.293
0.352
0.289
0.353
0.270
0.336
0.296
0.348
0.257
0.325
0.274
0.337
0.297
0.349
0.263
0.332
0.340
0.374
0.283
0.342
0.294
0.330
192
0.338
0.375
0.340
0.378
3.312
1.384
0.342
0.379
0.317
0.402
0.339
0.385
0.359
0.395
0.383
0.418
0.332
0.380
0.376
0.396
0.314
0.369
0.314
0.382
0.380
0.400
0.320
0.374
0.402
0.414
0.340
0.379
0.365
0.375
336
0.367
0.401
0.373
0.403
3.291
1.388
0.346
0.398
0.332
0.396
0.361
0.406
0.392
0.423
0.448
0.465
0.360
0.407
0.424
0.431
0.316 0.381
0.329
0.384
0.428
0.432
0.313
0.376
0.452
0.452
0.366
0.400
0.376
0.390
720
0.388
0.424
0.398
0.430
3.257
1.357
0.392
0.415
0.342
0.408
0.445
0.470
0.425
0.451
0.605
0.551
0.419
0.451
0.426
0.444
0.388
0.427
0.379
0.422
0.427
0.445
0.392
0.433
0.462
0.468
0.397
0.431
0.416
0.433
Avg
0.340
0.382
0.346
0.386
3.095
1.352
0.339
0.380
0.316
0.384
0.355
0.401
0.367
0.405
0.431
0.447
0.345
0.394
0.381
0.405
0.318
0.375
0.324
0.381
0.383
0.407
0.322
0.379
0.414
0.427
0.347
0.388
0.362
0.382
ETTm1
96
0.275
0.328
0.286
0.335
0.863
0.664
0.310
0.334
0.291
0.340
0.285
0.339
0.297
0.351
0.299
0.343
0.306
0.349
0.333
0.368
0.293
0.351
0.293
0.346
0.334
0.368
0.292
0.346
0.338
0.375
0.317
0.356
0.363
0.356
192
0.319
0.354
0.329
0.361
1.113
0.776
0.348
0.362
0.327
0.365
0.327
0.365
0.338
0.377
0.335
0.365
0.335
0.366
0.376
0.390
0.329
0.362
0.333
0.370
0.377
0.391
0.332
0.368
0.374
0.387
0.358
0.381
0.388
0.375
336
0.353
0.374
0.358
0.379
1.267
0.832
0.376
0.391
0.360
0.381
0.356
0.382
0.374
0.400
0.369
0.386
0.364
0.384
0.408
0.413
0.352
0.383
0.369
0.392
0.426
0.420
0.365
0.391
0.410
0.411
0.386
0.401
0.416
0.392
720
0.409
0.407
0.416
0.411
1.324
0.858
0.440
0.423
0.415
0.417
0.419
0.414
0.431
0.425
0.425
0.421
0.413
0.413
0.475
0.448
0.408
0.412
0.416
0.420
0.491
0.459
0.416
0.417
0.478
0.450
0.430
0.431
0.460
0.418
Avg
0.339
0.366
0.347
0.372
1.142
0.782
0.369
0.378
0.348
0.375
0.347
0.375
0.360
0.388
0.357
0.379
0.355
0.378
0.398
0.405
0.345
0.377
0.353
0.382
0.407
0.410
0.351
0.381
0.400
0.406
0.373
0.392
0.406
0.385
ETTm2
96
0.157
0.244
0.164
0.250
2.041
1.073
0.170
0.245
0.164
0.254
0.163
0.252
0.176
0.265
0.167
0.260
0.161
0.251
0.179
0.263
0.168
0.261
0.166
0.256
0.180
0.264
0.166
0.256
0.187
0.267
0.189
0.277
0.205
0.273
192
0.213
0.285
0.218
0.288
2.249
1.112
0.229
0.291
0.223
0.295
0.216
0.290
0.231
0.305
0.224
0.303
0.215
0.289
0.250
0.309
0.215
0.289
0.223
0.296
0.250
0.309
0.222
0.293
0.249
0.309
0.241
0.315
0.275
0.316
336
0.269
0.322
0.271
0.322
2.568
1.238
0.303
0.343
0.279
0.330
0.268 0.324
0.282
0.338
0.281
0.342
0.267
0.326
0.312
0.349
0.278
0.337
0.274
0.329
0.311
0.348
0.272
0.324
0.321
0.351
0.286
0.348
0.329
0.350
720
0.351
0.377
0.361
0.380
2.720
1.287
0.373
0.399
0.359
0.383
0.420
0.422
0.361
0.388
0.397
0.421
0.352
0.383
0.411
0.406
0.341
0.378
0.362
0.385
0.412
0.407
0.351
0.381
0.408
0.403
0.375
0.402
0.437
0.411
Avg
0.248
0.307
0.254
0.310
2.395
1.177
0.269
0.320
0.256
0.315
0.267
0.322
0.263
0.324
0.267
0.332
0.249
0.312
0.288
0.332
0.250
0.316
0.256
0.317
0.288
0.332
0.253
0.314
0.291
0.333
0.273
0.336
0.311
0.337
Wins
11
16
0
2
0
0
0
2
2
2
0
0
1
0
0
0
5
1
0
0
8 3
1
0
0
0
2
3
0
0
0
1
0
0
    Zero-shot forecasting.
  • $\dagger$
    Taken from Wu et al. [2022a].
  • $\ddagger$
    Traffic/PEMS are often used during pre-training [Liu et al., 2025]. Thus, no zero-shot results are available.
Tab. 6 shows the full results for long-term forecasting. This work involves conducting all experiments three times using seeds 2021, 2022, and 2023, following the setup of prior research [Wu et al., 2021, Nie et al., 2023, Wang et al., 2024a]. We therefore present the standard deviation of our model and the second-best models in terms of MSE and MAE in Tab. 7. This table, along with our experiments described in 4(a) and 4(b), further underscores the robustness of xLSTM-Mixer.
Table 7: xLSTM-Mixer provides state-of-the-art performance at low variance across datasets. This table shows the average performance and average standard deviation over all four prediction lengths in Tab. 6. They are contextualized by the competitive baselines TimeMixer and TiDE.
Model xLSTM-Mixer (ours) TimeMixer TiDE
Metric MSE MAE MSE MAE MSE MAE
Weather 0.219
$\pm$0.000
0.250
$\pm$0.000
0.240
$\pm$0.010
0.271
$\pm$0.009
0.236
$\pm$0.001
0.282
$\pm$0.001
Electricity 0.153
$\pm$0.001
0.245
$\pm$0.001
0.182
$\pm$0.017
0.272
$\pm$0.006
0.159
$\pm$0.002
0.257
$\pm$0.001
Traffic 0.392
$\pm$0.000
0.253
$\pm$0.000
0.484
$\pm$0.015
0.297
$\pm$0.013
0.356
$\pm$0.001
0.261
$\pm$0.001
ETTh1 0.397
$\pm$0.001
0.420
$\pm$0.001
0.047
$\pm$0.002
0.440
$\pm$0.005
0.419
$\pm$0.000
0.430
$\pm$0.000
ETTh2 0.340
$\pm$0.001
0.382
$\pm$0.000
0.364
$\pm$0.008
0.375
$\pm$0.010
0.345
$\pm$0.002
0.394
$\pm$0.001
ETTm1 0.339
$\pm$0.000
0.366
$\pm$0.000
0.381
$\pm$0.003
0.395
$\pm$0.006
0.355
$\pm$0.000
0.378
$\pm$0.000
ETTm2 0.248
$\pm$0.001
0.307
$\pm$0.001
0.275
$\pm$0.001
0.323
$\pm$0.003
0.249
$\pm$0.000
0.312
$\pm$0.000
Figure 5: xLSTM-Mixer is statistically significantly better than all baselines except xLSTMTime. Shown is the critical difference diagram for the MSE at a horizon of $H=96$. Horizontal bars connect methods that are not significantly different at $p=0.05$.
Appendix F Full Results on GIFT-Eval
This section completes Tab. 2 by reporting the full GIFT-Eval results: the overall leaderboard in Tab. 8, the multivariate and univariate subsets in Tab. 9 and Tab. 10, and the domain-wise Top-10 in Tab. 11. We follow the official GIFT-Eval protocol and scoring. The tables report the aggregate CRPS and MASE, and ranks are computed by the CRPS aggregate (lower is better). Values reflect the online leaderboard at the time of submission.
Table 8: GIFT-Eval Overall Leaderboard (full). Lower is better.
Model MASE $\downarrow$ CRPS $\downarrow$ Rank (CRPS) $\downarrow$
TiRex $0.724$ $0.498$ 11
xLSTM-Mixer (ours) $0.780$ $0.510$ 22
TEMPO_ensemble $0.862$ $0.514$ 33
Toto_Open_Base_1.0 $0.750$ $0.517$ 44
TabPFN-TS $0.771$ $0.544$ 55
YingLong_300m $0.798$ $0.548$ 66
timesfm_2_0_500m $0.758$ $0.550$ 77
YingLong_110m $0.809$ $0.557$ 88
sundial_base_128m $0.750$ $0.559$ 99
YingLong_50m $0.822$ $0.567$ $10$
chronos_bolt_base $0.808$ $0.574$ $11$
chronos_bolt_small $0.822$ $0.577$ $12$
TTM-R2-Finetuned $0.756$ $0.583$ $13$
PatchTST $0.849$ $0.587$ $14$
Moirai_large $0.875$ $0.599$ $15$
TFT $0.915$ $0.605$ $16$
YingLong_6m $0.880$ $0.609$ $17$
Moirai_base $0.901$ $0.610$ $18$
iTransformer $0.893$ $0.620$ 1919
Chronos_large $0.870$ $0.647$ 2020
Moirai_small $0.946$ $0.650$ 2121
Chronos_base $0.876$ $0.652$ $22$
Chronos_small $0.892$ $0.663$ $23$
TimesFM $1.077$ $0.680$ $24$
VisionTS $0.863$ $0.755$ $25$
TIDE $1.091$ $0.772$ $26$
N-BEATS $0.938$ $0.816$ $27$
DLinear $1.061$ $0.846$ $28$
DeepAR $1.343$ $0.853$ $29$
TTM-R2-Zeroshot $1.020$ $0.873$ $30$
Lag-Llama $1.228$ $0.880$ $31$
TTM-R1-Zeroshot $1.079$ $0.891$ $32$
Auto_Arima $1.074$ $0.912$ $33$
Timer $1.136$ $0.970$ $34$
seasonal_naive $1.000$ $1.000$ $35$
Auto_Theta $1.090$ $1.244$ $36$
Naive $1.270$ $1.591$ $37$
Crossformer $2.574$ $1.637$ $38$
Auto_ETS $1.212$ $7.489$ $39$
Across all datasets, xLSTM-Mixer ranks 2 out of 39 models with CRPS 0.512 and MASE 0.780 (Tab. 8) and is the strongest supervised method without pretraining. On the univariate and multivariate subsets, xLSTM-Mixer also ranks 2 with CRPS 0.543 and 0.474, respectively (Tab. 9, Tab. 10). These results are consistent with the other evaluations and show that the method transfers well to probabilistic forecasting with a simple quantile head. In particular, because the multivariate CRPS is lower than the univariate CRPS (0.474 vs. 0.543), this further confirms our previous observations and design decision that mixing of variates is beneficial, suggesting that xLSTM-Mixer effectively shares information across related series (e.g., common seasonality or shocks) without domain-specific tuning in the quantile head.
The domain breakdown indicates robust behavior across heterogeneous sources. The Top-10 per-domain tables show a first place in Healthcare and typically top-five performance across the remaining domains (Tab. 11). This suggests that the model adapts well to different frequencies, scales, and noise levels without domain-specific architectural changes.
Table 9: GIFT-Eval Multivariate Leaderboard (full). Lower is better.
Model MASE $\downarrow$ CRPS $\downarrow$ Rank (CRPS) $\downarrow$
TEMPO_ensemble $0.752$ $0.427$ 11
xLSTM-Mixer (ours) $0.839$ $0.473$ 22
Toto_Open_Base_1.0 $0.767$ $0.484$ 33
TiRex $0.771$ $0.490$ 44
YingLong_300m $0.840$ $0.519$ 55
YingLong_110m $0.830$ $0.520$ 66
YingLong_50m $0.853$ $0.530$ 77
sundial_base_128m $0.778$ $0.530$ 88
TabPFN-TS $0.818$ $0.544$ 99
timesfm_2_0_500m $0.803$ $0.553$ $10$
PatchTST $0.906$ $0.556$ $11$
YingLong_6m $0.905$ $0.566$ $12$
iTransformer $0.940$ $0.589$ $13$
TTM-R2-Finetuned $0.808$ $0.591$ $14$
chronos_bolt_small $0.939$ $0.600$ $15$
chronos_bolt_base $0.925$ $0.604$ $16$
TFT $1.071$ $0.610$ $17$
Moirai_large $1.023$ $0.635$ $18$
Moirai_base $1.054$ $0.636$ 1919
Moirai_small $1.021$ $0.640$ 2020
Chronos_large $1.005$ $0.680$ 2121
Chronos_small $1.026$ $0.684$ $22$
Chronos_base $1.013$ $0.685$ $23$
TimesFM $1.497$ $0.717$ $24$
VisionTS $0.887$ $0.721$ $25$
N-BEATS $0.998$ $0.790$ $26$
Lag-Llama $1.221$ $0.798$ $27$
TIDE $1.284$ $0.812$ $28$
DLinear $1.228$ $0.817$ $29$
Crossformer $1.479$ $0.825$ $30$
TTM-R2-Zeroshot $1.072$ $0.833$ $31$
TTM-R1-Zeroshot $1.186$ $0.856$ $32$
Timer $1.141$ $0.883$ $33$
DeepAR $1.908$ $0.989$ $34$
seasonal_naive $1.000$ $1.000$ $35$
Auto_Arima $1.318$ $1.032$ $36$
Auto_Theta $1.022$ $1.141$ $37$
Naive $1.167$ $1.455$ $38$
Auto_ETS $1.346$ $4.996$ $39$
Table 10: GIFT-Eval Univariate Leaderboard (full). Lower is better.
Model MASE $\downarrow$ CRPS $\downarrow$ Rank (CRPS) $\downarrow$
TiRex $0.688$ $0.505$ 11
xLSTM-Mixer (ours) $0.737$ $0.541$ 22
TabPFN-TS $0.735$ $0.544$ 33
Toto_Open_Base_1.0 $0.737$ $0.545$ 44
timesfm_2_0_500m $0.724$ $0.549$ 55
chronos_bolt_base $0.725$ $0.552$ 66
chronos_bolt_small $0.739$ $0.559$ 77
Moirai_large $0.773$ $0.572$ 88
YingLong_300m $0.766$ $0.573$ 99
TTM-R2-Finetuned $0.717$ $0.576$ $10$
sundial_base_128m $0.729$ $0.583$ $11$
Moirai_base $0.795$ $0.589$ $12$
YingLong_110m $0.793$ $0.589$ $13$
TEMPO_ensemble $0.960$ $0.596$ $14$
YingLong_50m $0.799$ $0.598$ $15$
TFT $0.808$ $0.601$ $16$
PatchTST $0.805$ $0.613$ $17$
Chronos_large $0.775$ $0.622$ $18$
Chronos_base $0.780$ $0.627$ 1919
YingLong_6m $0.861$ $0.646$ 2020
Chronos_small $0.797$ $0.647$ 2121
iTransformer $0.857$ $0.647$ $22$
TimesFM $0.829$ $0.652$ $23$
Moirai_small $0.890$ $0.659$ $24$
TIDE $0.959$ $0.741$ $25$
DeepAR $1.016$ $0.758$ $26$
VisionTS $0.845$ $0.783$ $27$
Auto_Arima $0.912$ $0.826$ $28$
N-BEATS $0.892$ $0.837$ $29$
DLinear $0.944$ $0.869$ $30$
TTM-R2-Zeroshot $0.980$ $0.907$ $31$
TTM-R1-Zeroshot $1.001$ $0.920$ $32$
Lag-Llama $1.233$ $0.952$ $33$
seasonal_naive $1.000$ $1.000$ $34$
Timer $1.131$ $1.047$ $35$
Auto_Theta $1.147$ $1.332$ $36$
Naive $1.358$ $1.709$ $37$
Crossformer $4.000$ $2.824$ $38$
Auto_ETS $1.115$ $10.337$ $39$
Table 11: GIFT-Eval Domain Leaderboards (Top-10 per domain). Lower is better.
Domain Model MASE $\downarrow$ CRPS $\downarrow$ Rank $\downarrow$
Econ/Fin timesfm_2_0_500m $0.640$ $0.580$ 11
TiRex $0.746$ $0.709$ 22
chronos_bolt_small $0.816$ $0.743$ 33
TimesFM $0.824$ $0.761$ 44
chronos_bolt_base $0.799$ $0.762$ 55
Moirai_large $0.845$ $0.778$ 66
TabPFN-TS $0.810$ $0.785$ 77
Chronos_base $0.783$ $0.798$ 88
xLSTM-Mixer (ours) $0.975$ $0.805$ 99
Chronos_large $0.782$ $0.806$ $10$
Energy TiRex $0.820$ $0.589$ 11
TEMPO_ensemble $1.063$ $0.613$ 22
YingLong_300m $0.870$ $0.627$ 33
Toto_Open_Base_1.0 $0.876$ $0.628$ 44
xLSTM-Mixer (ours) $0.881$ $0.633$ 55
chronos_bolt_base $0.846$ $0.640$ 66
TabPFN-TS $0.879$ $0.641$ 77
sundial_base_128m $0.839$ $0.645$ 88
YingLong_110m $0.907$ $0.651$ 99
chronos_bolt_small $0.864$ $0.656$ $10$
Healthcare xLSTM-Mixer (ours) $0.522$ $0.403$ 11
TabPFN-TS $0.576$ $0.450$ 22
TTM-R2-Finetuned $0.559$ $0.460$ 33
Toto_Open_Base_1.0 $0.625$ $0.467$ 44
Chronos_large $0.599$ $0.472$ 55
TiRex $0.628$ $0.473$ 66
timesfm_2_0_500m $0.597$ $0.481$ 77
Chronos_base $0.644$ $0.513$ 88
Chronos_small $0.607$ $0.525$ 99
chronos_bolt_small $0.671$ $0.541$ $10$
Nature TEMPO_ensemble $0.601$ $0.317$ 11
chronos_bolt_base $0.667$ $0.327$ 22
TiRex $0.686$ $0.328$ 33
Toto_Open_Base_1.0 $0.736$ $0.348$ 44
timesfm_2_0_500m $0.624$ $0.350$ 55
chronos_bolt_small $0.704$ $0.351$ 66
YingLong_300m $0.746$ $0.354$ 77
YingLong_110m $0.743$ $0.356$ 88
sundial_base_128m $0.703$ $0.361$ 99
YingLong_50m $0.748$ $0.365$ $10$
Domain Model MASE $\downarrow$ CRPS $\downarrow$ Rank $\downarrow$
Sales TiRex $0.682$ $0.415$ 11
TabPFN-TS $0.695$ $0.419$ 22
timesfm_2_0_500m $0.700$ $0.419$ 22
TimesFM $0.701$ $0.421$ 44
chronos_bolt_base $0.694$ $0.422$ 55
Moirai_base $0.695$ $0.424$ 66
chronos_bolt_small $0.696$ $0.424$ 66
Toto_Open_Base_1.0 $0.705$ $0.424$ 66
PatchTST $0.691$ $0.426$ 99
iTransformer $0.699$ $0.430$ $10$
Transport Moirai_large $0.601$ $0.451$ 11
TiRex $0.624$ $0.468$ 22
Toto_Open_Base_1.0 $0.632$ $0.477$ 33
Moirai_base $0.637$ $0.478$ 44
xLSTM-Mixer (ours) $0.635$ $0.487$ 55
TTM-R2-Finetuned $0.627$ $0.496$ 66
timesfm_2_0_500m $0.645$ $0.501$ 77
sundial_base_128m $0.634$ $0.504$ 88
YingLong_300m $0.666$ $0.504$ 88
TFT $0.679$ $0.514$ $10$
Web/CloudOps TEMPO_ensemble $0.585$ $0.387$ 11
xLSTM-Mixer (ours) $0.779$ $0.457$ 22
Toto_Open_Base_1.0 $0.694$ $0.500$ 33
TiRex $0.716$ $0.518$ 44
sundial_base_128m $0.695$ $0.552$ 55
PatchTST $0.780$ $0.553$ 66
TabPFN-TS $0.727$ $0.573$ 77
YingLong_110m $0.814$ $0.575$ 88
iTransformer $0.823$ $0.575$ 88
YingLong_300m $0.846$ $0.584$ $10$
Figure 6: xLSTM-Mixer is statistically significantly better than most baselines. Shown is the critical difference diagram for the MSE of the median. Horizontal bars connect methods that are not significantly different at $p=0.05$.
Appendix G Visualizing Initial Token Embeddings
Fig. 7 shows how the learned initial tokens $\bm{\eta}$ reflect common patterns found in the datasets. A row-by-row inspection of the figure reveals that similar patterns are learned per dataset, albeit at different scales. Specifically, these patterns repeat in proportion to the forecast horizon. The fact that essentially the same patterns are learned for each dataset across different initializations and horizons supports the conclusion that they are data-driven and meaningful to the domain.
Figure 7: Initial tokens capture dataset characteristics. The plot illustrates the learned tokens across multiple datasets and prediction lengths. The lookback length is set to 96 for all evaluations. For clarity and the high noise levels of the data, only a single seed is shown for ETTm1 and ETTh2.
Appendix H Learning Cross-Variate Patterns
As the sLSTM refinement blocks $\mathcal{S}$ process the variates recurrently, it is insightful to assess the extent to which inter-variate relationships are effectively captured. To this end, we adopt a perturbation-based approach to compute attributions, approximating Shapley Values through sampling. Hereby, we use a zero baseline and follow the horizon aggregation method proposed by Kraus et al. [2025], where the forecasts over the entire horizon are aggregated into a single scalar, which serves as the target for the attribution computation. We visualize these Shapley-based feature attribution scores, illustrating the degree to which each output variate of the xLSTM-Mixer depends on each input variate. Fig. 8 demonstrates the ability of xLSTM-Mixer to model cross-variate relationships effectively. Due to the design of the sLSTM refinement module, which strides over the variates, each variate can only be influenced by the ones preceding it. This restriction is reflected in the attribution scores, which appear exclusively in the lower-left triangle.
(a) Weather dataset.
(b) ETTh2 dataset.
(c) ETTm1 dataset.
Figure 8: xLSTM-Mixer effectively learns cross-variate patterns, as this feature attribution of each output to input variate on the Weather dataset demonstrates.
xLSTM-Mixer fixes one variate ordering to learn multivariate relationships efficiently. We investigate its impact by randomly permuting variate orders and comparing results with the baseline. Tab. 12 shows the results for four such permutations over four horizons on the Weather and ETTm1 datasets. We observe that the specific ordering does play some role in forecasting performance. However, the standard ordering provided by the dataset sources already permits highly effective forecasting.
Table 12: xLSTM-Mixer provides strong performance regardless of the variate permutations. All measurements are averaged over three initializations each.
From Dataset
Perm. #1
Perm. #2
Perm. #3
Perm. #4
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Weather
96
0.143
0.184
0.149
0.189
0.146
0.187
0.147
0.188
0.148
0.188
192
0.186
0.226
0.192
0.229
0.191
0.229
0.192
0.229
0.192
0.230
336
0.236
0.266
0.242
0.269
0.241
0.269
0.242
0.269
0.240
0.269
720
0.310
0.323
0.310
0.323
0.310
0.323
0.310
0.323
0.310
0.323
ETTm1
96
0.275
0.328
0.278
0.331
0.276
0.330
0.277
0.330
0.275
0.329
192
0.319
0.354
0.321
0.356
0.321
0.356
0.320
0.355
0.319
0.355
336
0.353
0.374
0.355
0.376
0.355
0.376
0.354
0.376
0.354
0.376
720
0.409
0.407
0.412
0.409
0.413
0.410
0.413
0.410
0.413
0.410
Electricity
96
0.126
0.218
0.127
0.220
0.126
0.218
0.127
0.219
0.125
0.218
192
0.144
0.235
0.145
0.237
0.144
0.235
0.145
0.236
0.144
0.235
336
0.157
0.250
0.160
0.252
0.159
0.251
0.157
0.248
0.159
0.250
720
0.183
0.276
0.230
0.315
0.225
0.312
0.206
0.295
0.218
0.306
Appendix I Ensembling over More than Two Views
Tab. 13 shows the results for ensembling varying numbers of views. Note that even if $E>2$, the first two views are always fixed to be forward and backward.
As we can see, the base version of using two views is the overall best choice. We observed this on multiple datasets and show ETTh2 as a representative sample.
Table 13: Ensembling over more than two views does not yield further benefits. Results are for different numbers of views $E$ for varying forecast horizons $H$ on ETTh2. The values in parentheses represent the relative change over $E=1$ (only forward view), where lower numbers are better. $E=2$ is the base variant of xLSTM-Mixer, showing the strongest improvement each.
$H=96$ $H=192$ $H=336$ $H=720$
$E$ MSE MAE MSE MAE MSE MAE MSE MAE
1 0.274 +0.00% 0.330 +0.00% 0.345 +0.00% 0.375 +0.00% 0.377 +0.00% 0.402 +0.00% 0.397 +0.00% 0.426 +0.00%
2 0.267
-2.55%
0.329
-0.30%
0.338
-2.03%
0.375
+0.00%
0.367
-2.65%
0.401
-0.25%
0.388
-2.27%
0.424
-0.47%
5 0.273
-0.36%
0.333
+0.91%
0.340
-1.45%
0.380
+1.33%
0.374
-0.80%
0.408
+1.49%
0.393
-1.01%
0.430
+0.94%
8 0.275
+0.36%
0.336
+1.82%
0.344
-0.29%
0.382
+1.87%
0.373
-1.06%
0.407
+1.24%
0.408
+2.77%
0.439
+3.05%
10 0.274
+0.00%
0.336
+1.82%
0.345
+0.00%
0.383
+2.13%
0.372
-1.33%
0.405
+0.75%
0.411
+3.53%
0.441
+3.52%
Appendix J Outlook: Classification
Tab. 14 compares xLSTM-Mixer to common time series classification models and standard benchmark datasets. Summarizing the individual results, xLSTM-Mixer is on par with the best model, ModernTCN [Donghao and Xue, 2023].
To adapt xLSTM-Mixer for classification, we replace the final regression projection layer with a single fully connected classification head followed by a softmax activation. We train this variant end-to-end on labeled sequences using cross-entropy loss, applying the same data preprocessing and augmentation pipeline as in our forecasting experiments.
These results highlight xLSTM-Mixer’s flexibility: with minimal architectural changes and no specialized classification tricks, it achieves state-of-the-art performance while preserving the same core building blocks used for forecasting. This unified design suggests that xLSTM-Mixer can serve as a general backbone for a wide range of sequence modeling tasks.
Table 14: xLSTM-Mixer is effective at time series classification. We report the averaged accuracy in percent. Adapted from Wu et al. [2022a].
Classical
Recurrent
SSM
Transformer
MLP
Convolutional
DTW
XGBoost
Rocket
xLSTM-
LSTM
LSTNet
S4
Trans.
Re.
In.
Pyra.
Auto.
Station.
FED.
ETS.
Flow.
iTransf.
DLin.
LightTS
TiDE
TimesNet
M.TCN
Datasets
[1994]
[2016]
[2020]
Mixer
[1997]
[2018]
[2022]
[2017]
[2020]
[2021]
[2021]
[2021]
[2022]
[2022]
[2022]
[2022b]
[2023]
[2023]
[2023]
[2023]
[2022a]
[2023]
EthanolConcentration
32.3
43.7
45.2
31.7
32.3
39.9
31.1
32.7
31.9
31.6
30.8
31.6
32.7
28.1
31.2
33.8
28.1
32.6
29.7
27.1
35.7
36.3
FaceDetection
52.9
63.3
64.7
68.9
57.7
65.7
66.7
67.3
68.6
67.0
65.7
68.4
68.0
66.0
66.3
67.6
66.3
68.0
67.5
65.3
68.6
70.8
Handwriting
28.6
15.8
58.8
31.8
15.2
25.8
24.6
32.0
27.4
32.8
29.4
36.7
31.6
28.0
32.5
33.8
24.2
27.0
26.1
23.2
32.1
30.6
Heartbeat
71.7
73.2
75.6
77.7
72.2
77.1
72.7
76.1
77.1
80.5
75.6
74.6
73.7
73.7
71.2
77.6
75.6
75.1
75.1
74.6
78.0
77.2
JapaneseVowels
94.9
86.5
96.2
97.5
79.7
98.1
98.4
98.7
97.8
98.9
98.4
96.2
99.2
98.4
95.9
98.9
96.6
96.2
96.2
95.6
98.4
98.8
PEMS-SF
71.1
98.3
75.1
91.5
39.9
86.7
86.1
82.1
82.7
81.5
83.2
82.7
87.3
80.9
86.0
83.8
87.9
75.1
88.4
86.9
89.6
89.1
SelfRegulationSCP1
77.7
84.6
90.8
93.6
68.9
84.0
90.8
92.2
90.4
90.1
88.1
84.0
89.4
88.7
89.6
92.5
90.2
87.3
89.8
89.2
91.8
93.4
SelfRegulationSCP2
53.9
48.9
53.3
59.8
46.6
52.8
52.2
53.9
56.7
53.3
53.3
50.6
57.2
54.4
55.0
56.1
54.4
50.5
51.1
53.4
57.2
60.3
SpokenArabicDigits
96.3
69.6
71.2
99.3
31.9
100.0
100.0
98.4
97.0
100.0
99.6
100.0
100.0
100.0
100.0
98.8
96.0
81.4
100.0
95.0
99.0
98.7
UWaveGestureLibrary
90.3
75.9
94.4
89.6
41.2
87.8
85.9
85.6
85.6
85.6
83.4
85.9
87.5
85.3
85.0
86.6
85.9
82.1
80.3
84.9
85.3
86.7
Average Accuracy
67.0
66.0
72.5
74.1
48.6
71.8
70.9
71.9
71.5
72.1
70.8
71.1
72.7
70.7
71.0
73.0
70.5
67.5
70.4
69.5
73.6
74.2